### v1

11institutetext: Institute for Astronomy, School of Physics, Zhejiang University, Hangzhou 310027, China
11email: [llh_astro; bbliu]@zju.edu.cn
22institutetext: Department of Physics, National University of Singapore, Singapore 117542, Singapore
22email: zekai77@u.nus.edu
Resonance Capture and Stability Analysis for Planet Pairs under Type I Disk Migration
Linghong Lin 11    Beibei Liu 11⋆⋆    Zekai Zheng 1122
We present a theoretical model to investigate a two-planet pair that undergoes convergent type I migration in a gaseous protoplanetary disk and traps into the first-order mean motion resonance. Our study identifies the conditions for resonant capture and explores the subsequent dynamical stability of the system. We derive the analytical criteria for planets with an arbitrary mass ratio and validate through numerical N-body simulations. Slow migration and weak eccentricity damping are required for resonant capture, the latter of which has not received enough attention in the literature. Once capture into resonance, their following stability can be classified into three regimes: stable trap, overstable trap and escape. Notably, resonant capture remains stable when the inner planet significantly outweighs the outer one. However, the subsequent evolution can be diverse when the mass of the inner planet is lower or comparable to that of the outer planet. Stability weakens as the relative strength between migration and eccentricity damping increases. The key results can be comprehensively demonstrated in a $\tau_{a}$-$\tau_{a}/\tau_{\rm e}$ plot, where $\tau_{a}$ and $\tau_{\rm e}$ are orbital decay and eccentricity damping timescales, respectively.
Key Words.:
planets and satellites: dynamical evolution and stability – celestial mechanics – methods: analytical
1 Introduction
Low-mass planets embedded in gaseous protoplanetary disks interact with disk gas and undergo type I orbital migration (Kley & Nelson, 2012; Paardekooper et al., 2023). Resonance capture can be a natural outcome of the planet pairs featured with convergent migration (Lee & Peale, 2002). The key question is to understand the under which conditions this capture can occur. The classical pendulum model pointed out that the trapping holds when the time planet migrates across the resonance width is longer than the time resonant angle librates (Murray & Dermott, 1999; Ogihara & Kobayashi, 2013; Batygin, 2015; Liu et al., 2015). A new criterion was derived by Huang & Ormel (2023), where they analyzed the resonant equilibrium state under the disk migration effects in a restrict three-body framework. Batygin & Petit (2023) provided an equivalent criterion using the Hamiltonian theory and extended it to the context of the unrestricted problem.
On the other hand, Goldreich & Schlichting (2014) first investigated the long-term stability of the resonant pair after capture in a gaseous disk. They found that under certain conditions, the resonant equilibrium state can become overstable, indicating that the planets may temporarily or permanently escape from resonance. Deck & Batygin (2015) conducted a similar analysis based on the Hamiltonian theory. Xu & Lai (2017) extended the framework to second-order mean motion resonances. Their study highlighted that the mass ratio of the planet pair is a crucial parameter in determining the overstability. Nesvorný et al. (2022) applied the overstable libration to account for the high eccentricity and large amplitude of the resonant angle observed in TOI-216b. Recently, Afkanpour et al. (2024) verified this overstable resonant trap using hydrodynamical simulations. They found that the outcome is related to the disk properties, such as disk mass and viscosity.
In this paper, we develop a unified framework for a two-planet pair that undergoes convergent type I migration and gets trapped at the first-order mean motion resonance ($j$:$j-1$ MMR). We investigate the conditions under which planets can be captured in resonances and examine the following stability properties. The key analytical criteria are derived for planets with arbitrary mass ratios and the outcomes are summarized in a $\tau_{\rm a}{-}\tau_{\rm a}/\tau_{\rm e}$ diagram.
The paper is organized as follows. We present an analytical model accounting for the dynamics of planets at resonances in Sec. 2. In Sec. 3 we perform N-body simulations and compare the results with the analytical predictions. In Sec. 4 we discuss several implications as well as potential caveats in our model. Finally, we summarize the key results in Sec. 5.
2 Dynamics of resonance trapping
In this section, we discuss the dynamical evolution of a two-planet pair in a gaseous protoplanetary disk after their convergent type-I migration. We consider two planets of masses $m_{\rm i}$ and $m_{\rm o}$ orbiting a central star of $M_{\star}$ on coplanar orbits. Their reduced masses and the mass ratio between the inner and outer ones are $\mu_{\rm i,o}{=}m_{\rm i,o}/M_{\rm\star}$ and $q{=}m_{\rm i}/m_{\rm o}$, respectively. For notations the subscripts ‘i’ and ‘o’ refer to the corresponding quantities of inner and outer planets.
We present the analytical criteria of resonance trapping and stability in Sec. 2.1 and Sec. 2.2, respectively. A unified picture is summarized in Sec. 2.3.
2.1 Resonance trapping criteria
The dissipative effect of planet-disk interaction can be parameterized by (Teyssandier & Terquem, 2014)
$\frac{1}{a}\frac{da}{dt}$ $=$ $-\frac{1}{\tau_{a}}-\frac{2e^{2}}{\tau_{e}},$ (1a)
$\frac{1}{e}\frac{de}{dt}$ $=$ $-\frac{1}{\tau_{e}},$ (1b)
where $\tau_{a}$ and $\tau_{e}$ are the orbital decay and eccentricity damping timescales, respectively. The orbital elements $a$, $n$, $e$, $\lambda$, $\varpi$, denote the semimajor axis, mean motion, eccentricity, mean longitude and longitude of pericenter of the planet.
Figure 1: Resonance crossing of a two-planet system with convergent migration. The plots provide the time evolution of the outer-to-inner planet period ratio (panel a), the two planets’ eccentricities (panel b), resonant angle (panel c) and phase curve (panel d). The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, the planet masses are $m_{\rm i}=1\ M_{\oplus}$, $m_{\rm o}=10\ M_{\oplus}$, and the migration parameters are $\tau_{a}=1.1\times 10^{5}yr$ and $\tau_{a}/\tau_{e}=1.5\times 10^{3}$. The planets directly cross the $2$:$1$ resonance at $t{\simeq}6\times 10^{3}$ yr and their eccentricities simultaneously excite to $10^{-2}$ and rapidly decline afterward. The resonant angle of the inner planet jumps from the initial values to $\pi$ during the crossing and oscillates over $2\pi$ afterwards. In phase space, the inner planet is not captured by the non-zero resonant fixed point and ultimately evolves towards the origin.
For the $j$:$j-1$ order mean motion resonances (MMRs), the resonant angles and the time-averaged disturbing functions for the inner and outer planets are given by (Murray & Dermott, 1999) $\varphi_{\rm i}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm i}$, $\varphi_{\rm o}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm o}$, and $R_{\rm i}{=}Gm_{\rm o}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+\alpha R_{\rm E})$, $R_{\rm o}{=}Gm_{\rm i}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+R_{\rm I}/\alpha^{2})$, where $\alpha{=}a_{\rm i}/a_{\rm o}$ is the two planets’ semi-major axis ratio, $R_{\rm D}^{\rm sec}$ and $R_{\rm D}^{\rm res}$ are the secular and resonant terms, $R_{\rm E}$ and $R_{\rm I}$ are the indirect part due to an external and internal perturber.
We assume small eccentricities of the planets, and expand the disturbing functions to the first order in $e_{\rm i}$ and $e_{\rm o}$, which gives $R_{\rm D}^{\rm sec}{=}0$, $R_{\rm D}^{\rm res}{=}e_{\rm i}f_{\rm d,i}\cos{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}\cos{\varphi_{\rm o}}$,$R_{\rm E}{=}-2e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2}$, and $R_{\rm I}{=}-\frac{1}{2}e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2}$, where $\delta_{\rm j,2}$ is the usual Kronecker symbol. Detailed expressions of $f_{\rm d,i}$ and $f_{\rm d,o}$ can be found in Table 8.1 of Murray & Dermott (1999).
Taking orbital decay and eccentricity damping into account, Lagrange’s equations can be written as
$\dot{n_{\rm i}}$ $=-3(j-1)\mu_{\rm o}n_{\rm i}^{2}\alpha(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm i}}{2\tau_{a,\rm i}}+\frac{3n_{\rm i}e_{\rm i}^{2}}{\tau_{e,\rm i}},$ (2a)
$\dot{n_{\rm o}}$ $=3j\mu_{\rm i}n_{\rm o}^{2}(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm o}}{2\tau_{a,\rm o}}+\frac{3n_{\rm o}e_{\rm o}^{2}}{\tau_{e,\rm o}},$ (2b)
$\dot{e_{\rm i}}$ $=-\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}\sin{\varphi_{\rm i}}-\frac{e_{\rm i}}{\tau_{e,\rm i}},$ (2c)
$\dot{e_{\rm o}}$ $=-\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}}-\frac{e_{\rm o}}{\tau_{e,\rm o}},$ (2d)
$\dot{\varpi_{\rm i}}$ $=\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}\frac{\cos{\varphi_{\rm i}}}{e_{\rm i}},$ (2e)
$\dot{\varpi_{\rm o}}$ $=\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}}\frac{\cos{\varphi_{\rm o}}}{e_{\rm o}},$ (2f)
where $f_{\rm d,o}^{{}^{\prime}}{=}f_{\rm d,o}-2\alpha\delta_{j,2}$. For instance, $f_{\rm d,i}{=}{-}1.19$, $f_{\rm d,o}^{{}^{\prime}}{=}0.43$ at the $2$:$1$ MMR.
When the planets are in MMRs, their orbital parameters reach the equilibrium state such that $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$, $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$. By substituting $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$ into Laplace equation, Eqs. (2a) and (2b) can be combined as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})$ (3)
$=\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
where we assume that $\tau_{a}^{-1}{=}\tau_{a,\rm o}^{-1}-\tau_{a,\rm i}^{-1}$ is the relative orbital decay rate with respect to convergent migration, and $\tau_{e}{=}\tau_{e,\rm i}{=}\tau_{e,\rm o}$ represents the eccentricity damping time-scale. Substituting $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ into Eqs. (2c) and (2d) to eliminate $\sin{\varphi_{\rm i,o}}$ and $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$ into Eqs. (2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$, the equilibrium eccentricities can be derived from Eq. (3) such that
$e_{\rm i,eq}^{2}$ $=$ $\frac{\tau_{e}}{2\tau_{a}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4a)
$e_{\rm o,eq}^{2}$ $=$ $\frac{\tau_{e}}{2\tau_{a}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4b)
where $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$.
Our results are analogous to Terquem & Papaloizou (2019), except for $\tau_{a}$ and $\tau_{e}$ as defined. In the limit of the mass of inner planet much lower than the outer one ($q{\ll}1$), the eccentricity of inner planet in Eq. (4a) can be simplified as
$e_{\rm i,eq}=\sqrt{\frac{1}{2j}\frac{\tau_{e}}{\tau_{a}}},$ (5)
in agreement with Eq. (24) in Huang & Ormel (2023). We note that due to differing signs of $f_{\rm d,i}$ and $f_{\rm d,o}^{{}^{\prime}}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ requires that $\sin{\varphi_{\rm i}}{>}0$ and $\sin{\varphi_{\rm o}}{<}0$ for $j$:$j-1$ MMRs. Considering $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$, we can derive $\sin{\varphi_{\rm i}}{=}-\sin{\varphi_{\rm o}}$, which implies $\varphi_{\rm i}-\varphi_{\rm o}{=}\pm\pi$. This means that the conjunctions between planet pair in a mean motion resonance always occur at either pericentre or apocentre.
Notably, the right-hand side of Eq. (3) represents the planet-disk interaction that leads convergent orbital migration between two planets, while the left-hand side of Eq. (3) represents their resonant gravitational interaction, keeping them away from getting closer. When the effect of planet-disk interaction is strong enough to overcome the resonant interaction, the planet pair cannot stay in resonance. Strong planet-disk interaction can be further classified into two regimes: fast migration and strong eccentricity damping. In the first regime the exchange rate of semimajor axis $\dot{a}/a$ in Eq. (3) is driven by the a-decay term ($\tau_{a}^{-1}{\gg}e^{2}\tau_{e}^{-1}$). In the second regime, however, the contribution of the e-damping term becomes comparable to that of the a-decay term. We will derive the resonance trapping conditions in the following subsections.
Figure 2: Resonance crossing of a two-planet system with convergent migration. The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, and the migration parameters are $\tau_{a}=10^{5}yr$ and $\tau_{a}/\tau_{e}=6\times 10^{2}$. The planets are trapped in the $2:1$ resonance at $t{\simeq}10^{4}$ yr and their eccentricities simultaneously excite to $2\times 10^{-2}$, which remain stable thereafter. The resonant angle of the inner planet is excited when the period ratio evolves to 2:1, subsequently oscillating and gradually converging to the resonant equilibrium value, which it maintains. In phase space, the inner planet is captured by the non-zero resonant fixed point and subsequently evolves in conjunction with this fixed point, reflecting the static equilibrium state of the resonance..
2.1.1 Slow migration
When the planets undergo fast convergent migration, they directly bypass the resonance before the eccentricities reach to the equilibrium values. In this case $\dot{a}/a$ is driven by $\tau_{a}$ but not $\tau_{e}$, and Eq. (3) can be rewritten as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{2\tau_{a}},$ (6)
where the e-damping term is ignored. Here the eccentricities are purely excited by the resonant term, which can be deduced from the classical pendulum model (see appendix Sec. Appendix A: The derivation of the excited eccentricity in the classical pendulum model for derivation)
$e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}=-\frac{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}.$ (7)
Combing Eqs. (6) and (7), the resonance trapping criterion of slow migration can be obtained as
$\tau_{a}>\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}.$ (8)
We further simplify the above criterion as
$\begin{cases}{\displaystyle\tau_{a}>\frac{1}{2n_{\rm i}(3(j-1))^{1/3}}\frac{1}{(\mu_{\rm o}\alpha f_{\rm d,i})^{4/3}}\ \sim\mu_{\rm o}^{-4/3}}\ \ \hfill\mbox{ at $q\ll 1$},\\ {\displaystyle\tau_{a}>\frac{1}{2n_{\rm o}(3j)^{1/3}}\frac{1}{(\mu_{\rm i}f_{\rm d,o}^{{}^{\prime}})^{4/3}}\ \sim\mu_{\rm i}^{-4/3}}\ \ \ \hfill\mbox{at $q\gg 1$}.\\ \end{cases}$ (9)
This indicates that the threshold migration rate increases with the planet mass. Intuitively, the planet resonant interaction becomes stronger as their masses increase. Therefore, the planets require faster migration in order to break the resonance.
2.1.2 Weak eccentricity damping
When the planets migrate relatively slow, they have time to reach the equilibrium state. Substituting the equilibrium eccentricity into the right side of Eq. (3), it turns out that the magnitude of e-damping term is comparable to that of the a-decay term. In this circumstance, the e-damping term becomes a key factor for resonance trapping.
Considering the limit of $\sin{\varphi_{\rm i}}{=}1$ and $\sin{\varphi_{\rm o}}{=}-1$, Eq. (3) can be rewritten as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$ (10)
where the left-hand side of the planetary interaction maintains a finite magnitude at resonances (${\propto}\mu ne$). If eccentricity damping is very strong, the orbital circularization leads large $\dot{a}/a$,, and the resonance can no longer hold. One example of resonance crossing is shown in Figure 1. Hence, resonance trapping requires weak e-damping. Substituting Eq. (10) with equilibrium values of $e_{\rm i}$ and $e_{\rm o}$ from Eqs. (4a) and (4b), we obtain the resonance trapping criterion in the weak e-damping regime as
$\tau_{e}\tau_{a}>\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}.$ (11)
In the limit of $q{\ll}1$, the trapping criterion can be simplified as
$\tau_{e}\tau_{a}>\frac{1}{2j(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i})^{2}},$ (12)
which is in consistent with Eq. (25) of Huang & Ormel (2023).
Altogether, the resonance trapping requires both relatively weak eccentricity damping and slow migration. Taking Eqs. (11) and (8) together, we formulate the necessary condition for $\tau_{a}$ and $\tau_{e}$ as
$\begin{cases}{\displaystyle(\tau_{e}\tau_{a})_{\rm weak,e}=\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vspace{0.1cm}\\ {\displaystyle(\tau_{a})_{\rm slow,mig}=\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}].\end{cases}$ (13)
Figure 3: Resonance crossing of a two-planet system with convergent migration. The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, and the migration parameters are $\tau_{a}=4\times 10^{5}yr$ and $\tau_{a}/\tau_{e}=5\times 10^{2}$. The planets are trapped in the $2:1$ resonance at $t{\simeq}3\times 10^{4}$ yr and their eccentricities first excite to $2\times 10^{-2}$, and starts to exhibit significant oscillations around its equilibrium value. The resonant angle of the inner planet remains at the equilibrium value and begins to oscillate after some time. In phase space, the inner planet initially evolves to a non-zero eccentricity resonant fixed point and subsequently starts oscillating around it. Over time, the system settles into a limit cycle around the fixed point, reflecting the dynamic equilibrium state of the resonance.
2.2 Post-evolution after resonant trapping
Here we assume that the planet pairs have satisfied the resonance trapping conditions in Sec. LABEL:sec:riteria and study their following dynamical evolution. The goal is to explore under which configuration the resonance state is temporary or long-term stable.
Our approach follows the analysis outlined in Section 2 of Xu & Lai (2017) (also see Chapter 8.8 of Murray & Dermott 1999). We first focus on the situation that the mass of the inner planet is much lower than the outer one ($m_{\rm i}{\ll}m_{\rm o}$, $q{\ll}1$). In the absence of planet-disk interactions, the system’s Hamiltonian $H$ near the $j$:$j-1$ MMR can be expressed as
$-H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (14)
where
$\varphi$ $=$ $j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm i},$ (15a)
$\Phi$ $=$ $[\frac{3(j-1)^{2}}{f_{\rm d,i}\mu_{\rm o}\alpha_{\rm 0}}]^{2/3}\frac{1-\sqrt{1-e_{\rm i}^{2}}}{2}\sim\mu_{\rm o}^{2/3}e_{\rm i}^{2},$ (15b)
$\eta$ $=$ $[\frac{1}{\sqrt{3}(j-1)f_{\rm d,i}\mu_{\rm o}}]^{2/3}[(j-1)\alpha_{\rm 0}^{-2/3}-j\alpha_{\rm 0}^{5/6}].$ (15c)
The first two parameters are a pair of conjugate coordinate and momentum, where $\varphi$ is the resonant angle and $\Phi$ represents the angular momentum deficit (AMD). The third parameter, $\eta$ is the dimensionless that quantifies the deviation level from the exact resonance. The orbital period ratio of the two planets is the exact integer ratio at $\eta{=}0$. The initial semi-major axis ratio of the two planet $\alpha_{\rm 0}\equiv\alpha[1+(j-1)e_{\rm i}^{2}]$ is the conjugate momentum of a fast angle, which can be derived from Eq.8-92 of Murray & Dermott (1999). We note that in the absence of planet-disk interactions, the parameters $\eta$ and $\alpha_{0}$ remain constant.
As we will demonstrate later, the effects of planet-disk interaction ($\tau_{\rm a},\tau_{\rm e}$) determines the stability of the planet pair in resonance.
2.2.1 stability/overstability criterion
When accounting for the planet-disk interaction, the equations of motion of the system can be obtained from
$\dot{\Phi}{=}-\frac{\partial H}{\partial\varphi}+\frac{\partial\Phi}{\partial e}\frac{de}{dt},\ \ \ \dot{\varphi}{=}\frac{\partial H}{\partial\Phi}.$ (16)
By incorporating the orbital decay and eccentricity damping terms in Eq.(1), Eq. (16) can be derived as
$\dot{\Phi}$ $=$ $\Phi^{1/2}\sin{\varphi}-\frac{2}{\tau_{e}}\Phi,$ (17a)
$\dot{\varphi}$ $=$ $-\eta-2\Phi+\frac{1}{2\sqrt{\Phi}}\cos{\varphi},$ (17b)
$\dot{\eta}$ $=$ $\frac{\partial\eta}{\partial\alpha_{0}}\frac{d\alpha_{0}}{dt}{=}-\frac{\beta}{\tau_{a}}+\frac{2(p+2j-2)\Phi}{(j-1)\tau_{e}},$ (17c)
where
$\beta\equiv|\frac{\partial\eta}{\partial\ln\alpha_{\rm 0}}|=[\frac{1}{\sqrt{3}(j-1)f_{\rm d,i}\mu_{\rm o}}]^{2/3}\frac{3}{2}(j-1)^{5/9}j^{4/9},$ (18)
and $\Phi{=}\beta[(j-1)e_{\rm i}^{2}/2]$.
The system reaches an equilibrium state at resonances when $\dot{\Phi}{=}0$, $\dot{\varphi}{=}0$, $\dot{\eta}{=}0$. Thus, the resonant parameters, to the lowest order, can be written as
$\Phi_{\rm eq}$ $=$ $\frac{(j-1)\beta\tau_{e}}{4j\tau_{a}},$ (19a)
$\varphi_{\rm eq}$ $\simeq$ $\frac{2}{\tau_{e}}\Phi_{\rm eq}^{1/2},$ (19b)
$\eta_{\rm eq}$ $\simeq$ $\frac{1}{2\sqrt{\Phi_{\rm eq}}}-2\Phi_{\rm eq}.$ (19c)
We then use the linear perturbation method to calculate these quantities ($\Phi{=}\Phi_{\rm eq}+\delta\Phi$ etc.). The corresponding linearized equations read
${\begin{bmatrix}\delta\dot{\Phi}\\ \delta\dot{\varphi}\\ \delta\dot{\eta}\end{bmatrix}=\begin{bmatrix}-\tau_{e}^{-1}&\Phi_{\rm eq}^{1/2}&0\\ -(0.25\Phi^{-3/2}+2)&-\tau_{e}^{-1}&-1\\ 4j\tau_{\rm e}^{-1}(j-1)^{-1}&0&0\end{bmatrix}\begin{bmatrix}\delta{\Phi}\\ \delta{\varphi}\\ \delta{\eta}\end{bmatrix}}.$ (20)
We assume that the solutions have the following forms $\delta\Phi,\delta\varphi,\delta\eta\propto e^{\lambda t}$ and $\lambda$ is a complex number. The characteristic equation can be written as
$\lambda^{3}+\frac{2}{\tau_{e}}\lambda^{2}+\left(2\Phi_{\rm eq}^{1/2}+\frac{1}{4\Phi_{\rm eq}}\right)\lambda+\frac{\beta}{\tau_{a}\Phi_{\rm eq}^{1/2}}=0.$ (21)
We provide a general approach to solve the real part solutions of Eq. (21) as folows. Suppose that a cubic equation has a form of $A\lambda^{3}+B\lambda^{2}+C\lambda+D{=}0$ where $A,B,C,D{>}0$. Assuming that $\lambda{=}\lambda_{\rm r}\pm i\lambda_{\rm i}$, it can be proved that when $D{<}BC/A$, the real part of eigenvalues $\lambda_{r}{<}0$. This consequently means that the trapping is stable where the trajectory of Hamiltonian $H$ converges to the fixed point (see an illustration in Figure 2). Substituting $A,B,C,D$ with Eq. (21), we obtain our stability criterion as
$\frac{\tau_{a}}{\tau_{e}}>\left(\frac{3}{f_{\rm d,i}}\right)^{2/3}\frac{(j-1)^{2/9}j^{4/9}}{2j}\mu_{\rm o}^{-2/3}\sim\mu_{\rm o}^{-2/3}.$ (22)
Our result is consistent with Eq. (26) of Goldreich & Schlichting (2014), where they constructed different forms of conserved quantities for derivations (see their Eq. 9).
Our analysis can also be extended to the case when the mass of the outer planet is much lower than that of the inner one ($m_{\rm i}{\gg}m_{\rm o}$, $q{\gg}1$). In such a circumstance, we find that the real part of eigenvalues of characteristic equation $\lambda_{\rm r}{<}0$ always holds. Importantly, this means that once the planet pair gets trapped in the resonance, the trapping is always stable as long as $m_{\rm i}{\gg}m_{\rm o}$.
Figure 4: Level curves of the Hamiltonian for different $\eta$, plotted in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$. The black (red) dots mark the stable (unstable) fixed points, the thick black line in panel (b) marks the separatrix. Each curve represents the conserved Hamiltonian in the absence of dissipitive disk forces. The color of the lines represents the values of the Hamiltonian. When considering the planet-disk interaction, during the convergent migration of the planet, H increases, which corresponds to a color shift from purple and green to yellow. When overstability occurs, the perturbations grow, causing the curve to expand, and H decreases, corresponding to a color shift from yellow and green to purple.
2.2.2 Escape criterion
When planets are in resonances and under the stable criterion, the resonant quantities ($\phi$,$\Phi$,$\eta$) remain fixed with a small amplitude liberation. Here we attempt to answer what occurs if the above stability criterion is not satisfied. We define such a state as overstablity. In this case, the perturbations grow exponentially over time (no longer in the linear regime). Whether the overstability leads planets directly escape or still stay in resonance can be further analyzed by the phase space topology.
Figure 4 shows the Hamiltonian curves for different $\eta$ values in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$. We note that $\eta$ represents the deviation from the exact resonance and the distance to the origin corresponds to the excited eccentricity strength ($\sqrt{X^{2}+Y^{2}}{=}\sqrt{\Phi}{\sim}e$).
Importantly, the value of $\eta$ sets the number of fixed points, and therefore determines the subsequent dynamical evolution. The fixed points can be obtained from $\partial H/\partial\varphi{=}\partial H/\partial\Phi{=}0$ (also see Eq.8-135 of Murray & Dermott 1999). It can be shown that, for $\eta{>}-3/2$, there is only one stable fixed point (black dot in Figure 4), whereas for $\eta{<}-3/2$, there are two stable fixed points and one unstable point (red dot in Figure 4).
In the overstable circumstances, the perturbation near the equilibrium grows significantly. When the equilibrium value $\eta_{\rm eq}{>}-3/2$, the system’s Hamiltonian curve expands as eccentricities excited during resonant capture. However, the trajectory still bounds to the only fixed point. As a result, the system’s eccentricity, period ratio, and resonant angle librate near the equilibrium. Following Goldreich & Schlichting (2014) and Xu & Lai (2017), we refer to this state as an overstable trap, which is also equivalent to circumstance described as the limit cycle in Afkanpour et al. (2024). One example of overstable trap can be found in Figure 3.
The separatrix occurs when $\eta_{\rm eq}{<}-3/2$ and the topology completely changes, as shown in Figure 4. In this case, the system’s Hamiltonian curve originally sits near the right stable fixed point. As the curve expands, it cross the separatrix and enter the circulation zone centered around the left stable fixed point (from the right to the left black dot in Figure 4). After then, the curve continuously shrinks towards the left black fixed point. As such, the eccentricity decreases, and the weaker resonant interaction cannot keep the planet pair in resonance anymore. The planets’ period ratio departs from the exact commensurability, and the system eventually escapes from the resonance. One example of escape is shown in Figure 5.
Based on Eq. (19c), $\eta_{\rm eq}{<}-3/2$ requires $\Phi_{\rm eq}{=}(j-1)\beta\tau_{\rm e}/(4j\tau_{a}){>}1$, which is corresponding to
$\frac{\tau_{a}}{\tau_{e}}<\left(\frac{3}{f_{\rm d,i}}\right)^{2/3}\frac{(j-1)^{8/9}j^{4/9}}{8j}\mu_{\rm o}^{-2/3}\sim\mu_{\rm o}^{-2/3}.$ (23)
The above escape criterion agrees with Eq. (28) of Goldreich & Schlichting (2014).
It is worth noting that this escape condition requires the system is already in the overstable regime. In the case where the mass of the outer planet is much lower than that of the inner one ($q\gg 1$), the resonance is always stable. As such, the escape does not occur even when $\eta_{\rm eq}{<}-3/2$.
Taking together, resonance trapping is always stable when the mass of the inner planet is much higher than that of the outer one ($q{\gg}1$). However, when the mass of the outer planet is much higher than that of the inner one ($q{\ll}1$), the outcome can be classified into three configurations: stable trap, overstable trap, or escape. We further extend the analysis to arbitrary mass ratios $q$ and derive the corresponding criteria as follows:
$\begin{cases}{\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}].\end{cases}$ (24)
where
$h(q)$ $=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (25)
$f(q)$ $=$ $1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (26)
Mention that $h(q)=\tau_{a}/\tau_{e}(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2})$, $f(q)=8(j-1)[1-(e_{\rm o,eq}/e_{\rm i,eq})^{2}]$. Eq. (24) reduces to Eqs. (22) and (23) in the limit $q\ll 1$. We show the detailed derivation in the Appendix B.
Figure 5: Resonance crossing of a two-planet system with convergent migration. The initial starting from the outer-to-inner planet period ratio is $2.2$ (slightly beyond the $2$:$1$ MMR), and migration parameters are $\tau_{a}=2.5\times 10^{5}$ and $\tau_{a}/\tau_{e}=10^{2}$. The period ratio between the inner and outer planets exhibits small oscillations upon entering resonance and ultimately decreases further away from the 2:1 ratio. The eccentricity of the inner planet undergoes significant oscillations around the resonant equilibrium value after resonance is established, eventually damping to a very small value. The resonant angle of the inner planet maintains large oscillations upon entering resonance and eventually diverges. In phase space, the inner planet initially oscillates around the resonant fixed point, but the oscillation range increases until it is ultimately captured by the origin.
Figure 6: The theoretical distribution of the four kinds of evolution near 2:1 MMR depending on different $\tau_{a}$ and $\tau_{a}/\tau_{e}$. When $\tau_{a}$ and $\tau_{a}/\tau_{e}$ are in the blue region, the two planets pass through the resonance because the eccentricity damping is too strong. When $\tau_{a}$ and $\tau_{a}/\tau_{e}$ are in the blown region, they can’t be trapped into the resonance since they migrate inward too fast. The green region indicates that the planets can maintain a stable resonance. The yellow area represents that although the resonance trapping is overstable, the planets can still maintain a 2:1 period ratio. The purple region indicates that the planets will eventually leave the 2:1 MMR position and migrate inward after overstability. The left panel shows the distribution when the outer planet is more massive than the inner one, and the right panel displays the distribution when the inner planet is more massive than the outer one.
2.3 A unified picture
We summarize different outcomes considering the dynamics evolution of planet pairs that undergo convergent migration. The illustrated runs from the numerical simulations in each regime are also presented.
  1. No trap (Figure 1): The system cannot reach the resonant equilibrium state and the planet pair directly bypasses the resonance. The planets’ eccentricities increase and their period ratio temporarily holds during the resonant passage. However, after rapid crossing the resonance, the eccentricity drops and the resonant angle varies from 00 to $2\pi$.
  2. Stable trap (Figure 2): The planets are locked in resonance with statically stable orbits. The planets’ period ratio, eccentricities, and resonant angle remain at the equilibrium values after capture. Meanwhile, the trajectory of the system converges towards a fix point in the phase space.
  3. Overstable trap (Figure 3): The planets are locked in resonance with dynamically stable orbits. Their eccentricities and resonance angle initially converge to the equilibrium values. After a while, the eccentricities and resonance angle liberate with increasing amplitudes. This state is nevertheless stable and the period ratio of planets remains constant after capture. In the phase space, the system’s trajectory initially approaches the fixed point, then begins to oscillate, and ultimately settles into a limit cycle around that point.
  4. Escape (Figure 5): The planets are locked in resonance with only temporarily stable orbits. Their period ratio, eccentricities and resonance angle initially converge to the equilibrium values. The planets’ eccentricities and resonance angle quickly oscillate with increasing amplitudes. This state becomes unstable and the planets eventually leave resonance. As such, the planets’ eccentricities decrease, and the period ratio continues to drop, and the resonant angle changes from liberation to circulation. In the phase space, the system’s trajectory initially approaches the fixed point; however, the oscillation gradually grows until it intersects the separatrix. The trajectory ultimately settles towards the coordinate origin.
We point out that although several literature studies have explored this issue in depth (Huang & Ormel, 2023; Goldreich & Schlichting, 2014; Deck & Batygin, 2015; Batygin, 2015), none have formulated a comprehensive model. Here we demonstrate a unified picture in Figure 6. The outcome is given in the $\tau_{a}$-$\tau_{e}$ parameter space. We describe how $\tau_{\rm a}$ and $\tau_{\rm e}$ influence the outcomes as follows.
If convergent migration is too rapid, the planets directly bypass the resonance without reaching the equilibrium state. As migration slows down, the effect of eccentricity damping becomes significant. Strong e-damping has two consequences: the planet resonant interaction gets weaken and semimajor axis decays due to orbital circularization. When the orbital decay surpasses the resonant term, the planets will no longer remain in resonance. In this regard, planets cannot get capture when $\tau_{\rm e}$ and/or $\tau_{\rm a}$ are short. No trap occurs in the upper and left part of Figure 6.
When planets are trapped in resonance, the stability is related to the degree of eccentricity excitation. The damping effect ($\tau_{\rm e}$) reduces the planet’s eccentricity, leading the system stable. On the other hand, convergent migration ($\tau_{\rm a}$) brings planets closer together, in turn exciting their eccentricities. Consequently, the relative eccentricity damping $\tau_{\rm a}/\tau_{\rm e}$ is the key parameter affecting the stability. As can be seen from top to bottom of Figure 6, system tends to be more unstable as $\tau_{\rm a}/\tau_{\rm e}$ decreases.
For the small inner planet case shown in Figure 6, if the relative damping $\tau_{\rm a}/\tau_{\rm e}$ is sufficiently strong, the resonant trapping is statically stable, and the final state of the system stays at a fixed point in the phase space (stable trap regime). As the relative damping weakens (Eq. 22), the fixed point becomes unstable, and the system’s trajectory begins to oscillate around the fixed point. It ultimately settles into a stable limit cycle (overstable trap). The limit cycle expands as $\tau_{\rm a}/\tau_{\rm e}$ continues to decrease. When the trajectory crosses the separatrix (Eq. 23), the system becomes dynamically unstable. In this regime, the oscillations grow over time and planets eventually leave resonance (escape). In contrast, for the case of a small outer planet shown in Figure 6, all traps are stable.
Figure 7: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q$ of $0.1$. Blue points indicate cases where the planets are not trapped by the resonance, green points indicate stable resonance capture, yellow points indicate overstable resonance capture, and purple points indicate that the planets eventually escape from the resonance. The simulations generally align with the schematic shown in Figure 6, except for the region near $\tau_{a}\sim 10^{5}\ \mathrm{yr}$ and $\tau_{a}/\tau_{e}\sim 3\times 10^{2}$. While the analytical model predicts overstable trapping in this regime, the simulations show escape from resonance. We hypothesize that as migration becomes faster, the perturbations grow stronger and become nonlinear, leading to amplified oscillation amplitudes and an increased likelihood of escaping from resonance.
Figure 8: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q$ of $10$. The simulations generally align with the schematic shown in Figure 6, except for the region near $\tau_{\rm a}{\sim}2{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm a}/\tau_{\rm e}{\sim}2\times 10^{2}$. While the analytical model predicts resonance trapping in this regime, the simulations show resonance crossing. We hypothesize that the discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect.
Figure 9: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q=1$. We compare our criteria with previous studies.We cite the trapping criterion from Huang & Ormel (2023) and the stable and escape criteria from Goldreich & Schlichting (2014). The solid lines represent our criteria, the dash lines represent the criteria cited from previous studies. In the case of equal-mass planets, it is evident that our criteria align more closely with the simulation results.
3 Simulations
In this section, we conduct numerical simulations for $2:1$ resonance to verify the different regimes and criteria analytically derived in the previous section. The numerical setup is described in Sec. 3.1. The comparison between simulations and analysis at two extreme mass ratios is presented in Sec. 3.2, and the extrapolation for any mass ratio is discussed in Sec. 3.3.
3.1 Numerical setup
We use the open-source N-body code REBOUND with the WHfast integrator (Rein & Liu, 2012). The simulations are confined in conplanar circumstances, and the accelerations for orbital decay and eccentricity damping are given by
${a_{\rm m}}=-\frac{\textbf{v}}{2\tau_{a}},\ \ {a_{\rm e}}=-2\frac{(\textbf{v}\cdot\textbf{r})\cdot\textbf{r}}{r^{2}\tau_{e}},$ (27)
which are implemented through REBOUNDx (Tamayo et al., 2020).
Initially, we place the inner and outer planets at $1$ and $1.7$ au, respectively, making them slightly away from the $2$:$1$ MMR. Both planets have circular orbits with random phase angles. To mimic convergent migration, we only let the outer planet undergo inward migration on a time-scale of $\tau_{\rm a}$, whereas the eccentricities of both planets are damped on a time-scale of $\tau_{\rm e}$. Simulations are terminated at $t{=}2\tau_{\rm a}$, and the final outcomes of the planets in each run are determined as described in Sec. 2.3.
We consider three cases for the masses of the planet-pair: small inner planet ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $q{=}0.1$), small outer planet ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, $q{=}10$), equal-mass planets ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, $q{=}1$). We vary different sets of $\tau_{\rm a}$ and $\tau_{\rm e}$ in each simulations.
3.2 Comparison at two extreme mass ratios
The dynamical outcomes of planet pairs that undergo convergent migration are shown in Figure 6. The results are dependent on the mass ratio of the planets and disk properties ($\tau_{\rm a}$ and $\tau_{\rm e}$). Basically four regimes are presented when the inner planet is lower than that of the outer one. The illustrated runs in each regime are provided in Figures 1,2,3 and 5.
Figure 8 shows the numerical results obtained from an illustrated case where the mass of inner planet is ten times lower than that of the outer one ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $q{=}0.1$). The numerical results shown in Figure 8 can directly be compared with the schematic presented in Figure 6.
We find that the simulations agree pretty well with the analytical estimates. No planets lead resonant trap (blue dots) in the upper left part of Figure 8, when the weak damping (upon the green line) and slow migration (left of the brown line) criteria break. When those trapping criteria are satisfied and as $\tau_{\rm a}/\tau_{\rm e}$ decreases, the simulations depict a transition from stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes. These results are indeed separated by the two analytical criteria proposed in Sect. 2 (yellow and purple lines).
However, there is one point in Figure 8 that simulations do not fully match the analysis, occurring at $\tau_{a}{\sim}10^{5}\ {\rm yr}$ and $\tau_{a}/\tau_{e}{\sim}3\times 10^{2}$. The theory predicts the planets in overstable trap regime whereas in simulations planets actually leave the resonance. We speculate that the reason is that as migration becomes faster, the perturbations intensify and become nonlinear, resulting in further increasing oscillation amplitudes around the equilibrium value. Consequently, the system becomes more likely to cross the separatrix and escape resonance.
Figure 8 demonstrates simulation results from the case when the mass of inner planet is ten times higher than that of the outer one ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, $q{=}10$), serving as a comparison to the schematic illustrated in Figure 6. Again, no trap (blue dots) occurs in the upper and left parts of Figure 8, when the weak damping and slow migration criteria are not satisfied. When these criteria are fulfilled, all traps correspond to stable trap (green dots).
We note that the no-trap regime occupies a large parameter space than theoretically expected. This is evident at $\tau_{\rm a}{\sim}2{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm a}/\tau_{\rm e}{\sim}2\times 10^{2}$, where the trapping criteria are fulfilled but the planets are not in resonance. This discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect (Eq. 6). When the eccentricity damping is considered, the interaction between the planets is further suppressed (lower $e$ in left-hand side of Eq. 3), leading to a longer migration timescale (right-hand side of Eq. 3). Consequently, in reality planets can bypass the resonance with slower migration rates compared to the simplified estimation from Eq. 6.
In conclusion, our analytical predictions align well with simulations in the circumstances of planet pairs with extreme mass ratios.
3.3 Comparison at arbitrary mass ratios
Notably, Huang & Ormel (2023) derived the weak damping trapping criterion at the small mass ratio limit of $q{\ll}1$. Goldreich & Schlichting (2014) investigated when the trap becomes overstability and/or escape in the same limit. In this work we derived the general criteria for trap and stability that are well suited to the arbitrary mass ratio (Eqs. 13 and 24). In this subsection we show the simulation comparison, whereas the detailed derivation of the $q$-dependent overstabity and escape criteria is provided in Appendix B.
Figure 9 illustrates the simulation results for the equal-mass planet case ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, $q{=}1$) and compares our criteria with those from Huang & Ormel (2023) and Goldreich & Schlichting (2014). The solid lines represent our analytical criteria, dashed lines correspond to criteria from previous studies, and colored dots depict the simulation results. We find that the new criteria clearly depict a better agreement with numerical results.
Figure 10: The plots illustrate the dependence of resonance trapping criteria on the semi-major axis (panel a) and the stellar mass (panel b) for a two-planet system, based on the minimum solar nebular model (MMSN). The planet masses are $m_{\rm i}=1\ M_{\oplus}$ and $m_{\rm o}=10\ M_{\oplus}$. In panel a, the stellar mass is fixed at $M_{\star}=1\ M_{\odot}$, while in panel b, the semi-major axis of the inner planet is fixed at $a_{\rm i}=1\ au$. The brown line represents the slow migration trapping criterion (Eq. 8), below which migration is too fast for resonance capture. Regions above the brown line indicate where capture is possible. The color gradient represents the strongest eccentricity damping $\tau_{a0}/\tau_{e0}$ that allows resonant capture (Eq. 11), with lighter colors meaning the planets can be in resonances with stronger eccentricity damping. In panel (a), as the semi-major axis increases, the critical value of $\tau_{a0}$ remains constant, while $\tau_{a0}/\tau_{e0}$ slightly increases. This suggests that planets are more easily trapped in resonances under stronger eccentricity damping as the semi-major axis increases. In panel (b), as the stellar mass decreases, the critical value of $\tau_{a0}$ decreases, and $\tau_{a0}/\tau_{e0}$ increases. This indicates that planets are more easily trapped in resonances under stronger eccentricity damping and faster migration as the stellar mass decreases.
4 Discussion
In this section, we connect our criteria to disk parameters in Sec. 4.1 and analyze how the criteria depend on the planet location and the mass of stellar host. Additionally, we investigate resonance trapping in the context of divergent migration in Sec. 4.2 and discuss several caveats of our work in Sec. 4.3.
4.1 Dependence on semi-major axis and stellar mass
Here we explore how the resonant trapping conditions depend on the radial distance and stellar mass. The detailed type I migration formulas for low-mass planets are given by $\tau_{a}^{-1}{\sim}\mu\mu_{\rm d}h^{-2}n$ and $\tau_{e}^{-1}\sim\mu\mu_{\rm d}h^{-4}n$, where $\mu_{\rm d}{=}\Sigma_{\rm d}a^{2}/M_{\rm*}$ is the mass ratio between disk and star, and $\Sigma_{\rm d}$ and $h$ are the disk surface density and aspect ratio, respectively.
Following the minimum mass solar nebula (MMSN, Hayashi 1981) model, the disk surface density and aspect ratio can be written as
$\Sigma_{\rm d}=$ $1700\left(\frac{a}{\rm 1\ au}\right)^{-3/2}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)\rm\ g/cm^{2},$ (28)
$h=$ $0.033\left(\frac{a}{\rm 1\ au}\right)^{1/4}\left(\frac{L_{\star}}{1\ L_{\odot}}\right)^{1/8}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{-1/2}$
$=$ $0.033\left(\frac{a}{\rm 1\ au}\right)^{1/4}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{-1/4},$ (29)
the latter of which $L_{\star}{\propto}M_{\star}^{2}$ is assumed for the pre-main sequence stars. By substituting $\Sigma_{\rm d}$ and $h$ into $\tau_{a}$ and $\tau_{e}$, we obtain their dependence on the semi-major axis and stellar mass as
$\tau_{a}={\tau_{a0}}\left(\frac{a}{\rm 1\ au}\right)^{3/2},\ \ \tau_{e}={\tau_{e0}}\left(\frac{a}{\rm 1\ au}\right)^{2}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{1/2},$ (30)
where $\tau_{a0}$ and $\tau_{e0}$ represent the values of $\tau_{a}$ and $\tau_{e}$ at $1$ au around a solar-mass host. By inserting $\tau_{a0}$ and $\tau_{e0}$ into Eq. (13), we can separate the dependence of resonance trapping on semi-major axis and stellar mass.
Figure 10 shows the a-dependency for resonance trapping, where $M_{\star}{=}1\ M_{\odot}$ and $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$. The region below the brown line (slow migration criterion) indicates that migration is too fast for resonance capture, while the region above suggests that capture is possible, which depends on $\tau_{a0}/\tau_{e0}$. The color in Figure 10 refers to the critical $\left(\tau_{a0}/\tau_{e0}\right)$ that allows resonant capture. A lighter color means the planets can be in resonances with higher $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ (stronger e-damping). As the semi-major axis increases, $(\tau_{a0})_{\rm slow,mig}$ remains constant and $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ slightly increases. This means that planets would be more easily trapped at resonances under stronger e-damping as $a$ increases. Nevertheless, the capture reveals an overall weak semi-major axis dependence.
Figure 10 shows $M_{\star}$-dependency for resonance trapping. We fix the inner planet at $1\ au$ and vary with the stellar mass. From the figure, as the stellar mass decreases, $(\tau_{a0})_{\rm slow,mig}$ decreases, and $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ increases. This implies that planets can be trapped under a wider range of $\tau_{a}$ and $\tau_{e}$ parameter space. In other words, planets are more likely to be capture into resonance as the stellar mass decreases.
Based on the MMSN model, we find that planets are more likely to get capture into resonance with increasing semi-major axis and/or decreasing stellar mass. However, we note that these results are considerably sensitive to the adopted disk model. The interpretation should be taken with caution.
4.2 Divergent migration
We have shown in the previous sections that the convergent planet migration coupled with eccentricity damping can result in resonance trapping. Literature theoretical work suggested that planets cannot be captured in resonance when the direction of their migration is divergent (Wong & Lee, 2024). Based on our analysis, here we also provide a simple argument that why resonance trapping cannot occur through divergent migration.
When the planet pair undergoes the divergent migration, the equilibrium state $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$ (if it exists) can be written as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})$ (31)
$=-\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
where the relative orbital decay term ($-\tau_{a}^{-1}$) is negative compared to Eq. (3). In order to derive equilibrium eccentricity, we repeat the procedure outlined in Sec. 2.1 by substituting $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ into Eqs. (2c) and (2d) to eliminate $\sin{\varphi_{\rm i,o}}$ and $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$ into Eqs. (2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$. Then Eq. (31) reads
$\frac{e_{\rm i}^{2}}{\tau_{\rm e}}(1+q\sqrt{\alpha})[j+(j-1)\frac{f_{\rm d,o}^{{}^{\prime}2}}{f_{\rm d,i}^{2}}q\sqrt{\alpha}]=-\frac{1}{2\tau_{a}}.$ (32)
Noticeably, the left-hand side of this equation is positive, whereas the right-hand side is negative. This indicates that there is no equilibrium solution in the divergent migration circumstance. We also conducted numerical simulations and consistently found that even a small amplitude of divergent migration is sufficient for planets to cross the resonance.
4.3 Caveats
Here discuss a few limitations in our model. First, for simplicity we treat $\tau_{\rm a}$ and $\tau_{\rm e}$ as free but constant parameters and two planets have the same eccentricity damping timescales. However, in reality $\tau_{\rm a}$ and $\tau_{\rm e}$ depend on the gas disk mass, planet mass and orbital distance. For instance, $\tau_{\rm a}$ and $\tau_{\rm e}$ decrease as planet mass increases, which scale with $\mu^{-1}$. The two trapping criteria exhibit steeper decreasing functions of $\mu$ such that $\tau_{\rm a,\ \text{slow\ mig}}{\propto}\mu^{-4/3}$, and $(\tau_{\rm a}\tau_{\rm e})_{\rm\text{weak\ damp}}{\propto}\mu^{-2}$. This suggests that planets become more easily trapped as their mass increases. On the other hand, since the planets can grow their masses at different rates, the planet mass ratio could be time varied. For instance, if the inner planet is initially more massive than the outer planet ($q{\gg}1$), the resonance trapping is the stable regime. When the outer planet grows and the mass ratio decreases, the system may transition into an overstable trap or even escape. Moreover, both hydrodynamic simulations and analytical work have shown that the migration and eccentricity damping rates differ from those in the near circular case when planets attain high eccentricities (Cresswell & Nelson, 2008; Ida et al., 2020). Future studies are required to investigate how these factors influence our analysis.
Second, we only consider a static disk. The configuration may change when disk dissipation is taken into account. In this case, the migration slows down as the gas density drops, making the resonance trapping easier. However, for stability/overstability and escape criteria, the situation is less straightforward. This is because $\tau_{\rm a}/\tau_{\rm e}$ depends on the aspect ratio of the disk but not the surface density. The disk thermal structure relies on the heating mechanism and therefore a specific disk model. As such, our framework cannot predict whether the planet pairs in the stable or overstable regimes will remain trapped as gas disk dissipates.
In addition, the disk is assumed to be laminar here. In a more realistic turbulent disk, density fluctuations generated by turbulence can result in additional stochastic torques (Goldberg & Batygin, 2023; Rein, 2012). This effect cannot be captured by our simplified approach that assumes static migration and eccentricity damping timescales. However, the outcomes of resonance trapping may be altered by considering the stochastic torques, which is a key focus of our future work.
5 Conclusion
We have built an analytical framework suited to study the resonant trapping, stability, and escape conditions for a two-planet pair that undergoes the convergent type I migration in a gaseous disk. The $2$:$1$ first order mean motion resonance is specifically investigated and the key outcomes are illustrated in Fig 6.
Planets can be trapped in resonances under the circumstances of slow migration and weak damping (Eqs. (8) and (11)). When the mass of inner planet is comparable or less massive than the outer planet ($q{\lesssim}1$), after capture the stability weakens as $\tau_{\rm a}/\tau_{\rm e}$ decreases. The system transitions from a stable trap to an overstable trap (Eq. 22), and finally to escape (Eq. 23). This corresponds to the resonant states changing from statically stable to dynamically stable, and ultimately to dynamically unstable. In contrast, when the mass of inner planet is much more massive than the outer planet ($q{\gg}1$), the traps remain permanently stable.
We provide the general analytical criteria for planets with an arbitrary mass ratio $q$ in Sec. 2. The corresponding formulas are also summarized as follows.
$\begin{cases}{\displaystyle(\tau_{e}\tau_{a})_{\rm weak,e}=\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vspace{0.1cm}\\ {\displaystyle(\tau_{a})_{\rm slow,mig}=\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}],\end{cases}$ (33)
where
$h(q)$ $=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (34)
$f(q)$ $=$ $1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (35)
We validate the theoretical predictions with numerical N-body simulations (Figures 8, 8 and 9). Furthermore, within the framework, we show that resonance trapping cannot occur when planets migrate divergently. We also explored the dependence of the semi-major axis and the stellar mass based on minimum mass solar nebular (MMSN) model, which indicates planets are more likely to be resonance-trapped with increasing semi-major axis and decreasing stellar mass.
Acknowledgements.
This work is supported by REBOUND (Rein & Liu, 2012) and REBOUNDX (Tamayo et al., 2020)
Appendix A: The derivation of the excited eccentricity in the classical pendulum model
When the planets are in resonance, the second time derivative of the resonant angle $\varphi_{\rm i}$ and $\varphi_{\rm o}$ can be written as
$\ddot{\varphi_{\rm i}}{=}\ddot{\varphi_{\rm o}}{=}j\dot{n_{\rm o}}+(1-j)\dot{n_{\rm i}},$ (A1)
where the contribution from $\dot{\varpi_{\rm i}}$ and $\dot{\varpi_{\rm o}}$ are neglected. With no additional effects from the gaseous disk such as orbital decay and eccentricity damping, Lagrange’s equations can be written as
$\dot{n_{\rm i}}$ $=-3(j-1)\mu_{\rm o}n_{\rm i}^{2}\alpha(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}}),$ (A2a)
$\dot{n_{\rm o}}$ $=3j\mu_{\rm i}n_{\rm o}^{2}(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}}),$ (A2b)
$\dot{e_{\rm i}}$ $=-\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}\sin{\varphi_{\rm i}},$ (A2c)
$\dot{e_{\rm o}}$ $=-\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}}.$ (A2d)
Substituting $\dot{n_{\rm o}}$ and $\dot{n_{\rm i}}$ in Eq. (A1) with Eqs. (A2a) and (A2b), and combing Eqs. (A2c) and (A2d), we obtain
$\ddot{\varphi}{=}3[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})\sin{\varphi},$ (A3)
$\dot{e_{\rm i}}f_{\rm d,i}-\dot{e_{\rm o}}f_{\rm d,o}^{{}^{\prime}}=-(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}})\sin{\varphi},$ (A4)
where we assume $\sin{\varphi}\equiv\sin{\varphi_{\rm i}}=-\sin{\varphi_{\rm o}}$. The equation for $\ddot{\varphi}$ can be simplified to the form
$\ddot{\varphi}{=}-\omega_{0}^{2}\sin{\varphi},$ (A5)
where we take
$\omega_{0}^{2}=-3[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}).$ (A6)
The solution can be described as a libration of $\varphi$. The total energy $E$ of the system is the sum of the kinetic energy and potential energy. Hence $E=\frac{1}{2}\dot{\varphi}^{2}+2\omega_{0}^{2}\sin^{2}{\varphi/2}$. It is clear that the maximum libration occurs when $\dot{\varphi}=0$ at $\varphi=\pm\pi$. This implies $E_{\rm max}=2\omega_{0}^{2}$. We now set the value of $E$ equal to $E_{\rm max}$ and consider $\dot{\varphi}$ given by
$\dot{\varphi}=\pm 2\omega_{0}\cos{\varphi/2}.$ (A7)
We can relate the variation of $\dot{\varphi}$ to the variation in $e_{\rm i,o}$ by means of Eq. (A4). This gives
$d(e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{d,o^{{}^{\prime}}}){=}\pm\frac{(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}})}{\omega_{0}}\sin{\frac{\varphi}{2}}d{\varphi}.$ (A8)
Substitute $\omega_{0}$ into Eq. (A8) and integrate it, we obtain
$e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}=-\frac{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}\cos{\frac{\varphi}{2}}.$ (A9)
The maximum eccentricity occurs when $\varphi=0$, which provides the eccentricity described in Sec. 2.1.1.
Appendix B: Extension of Stability and Escape Criteria to Arbitrary Mass Ratios
B.1: Escape Criterion
For the cases of $q\ll 1$ and $q\gg 1$, the system follows the same Hamiltonian equation:
$-H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (B1)
where
$\begin{cases}{\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}]^{2/3}\frac{e_{\rm i}^{2}}{4},\ \ \hfill[\mbox{$q\ll 1$}],\vspace{0.1cm}}\\ {\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm i}}\right)^{2/3}[\frac{j^{2}}{f_{\rm d,o}^{{}^{\prime}}}]^{2/3}\frac{e_{\rm o}^{2}}{4},\ \ \hfill[\mbox{$q\gg 1$}].}\end{cases}$ (B2)
Therefore, we hypothesize the existence of a form for $\Phi$ that satisfies the Hamiltonian equation and degenerates to Eq. (B3) in the extreme mass ratio cases:
$\Phi=\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}]^{2/3}\frac{e^{2}}{4},$ (B3)
where $e^{2}=e_{\rm i}^{2}+e_{\rm o}^{2}$. When $\eta_{\rm eq}{<}-3/2$ (equivalently $\Phi_{\rm eq}>1$), a separatrix point occurs and the system can cross this separatrix and escape from resonance (see Sec. 2.2.2). By substituting $e_{\rm i,eq}$ and $e_{\rm o,eq}$ (Eq. 4) into Eq. (B3), we obtain the general escape criterion as
$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}.$ (B4)
where
$h(q)$ $=$ $\tau_{a}/\tau_{e}\left(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2}\right)$ (B5)
$=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}}.$
B.2: Stability Criterion
The stability of the system decreases with the mass ratio $q$. When the outer planet is comparable or more massive than the inner planet ($q\lesssim 1$), the system may become overstable (see Eq. 22).However, all traps remain stable when the inner planet is more massive. Therefore, we introduce a factor $f(q)$ to fit the general stability criterion, which is given by
$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q),$ (B6)
where
$f(q)=1-\frac{e_{\rm o,eq}^{2}}{e_{\rm i,eq}^{2}}\ =1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (B7)
When $q\ll 1$, $f(q)\sim 1$, the system is dominated by the outer planet. As $q$ increases, $f(q)$ decreases, indicating that the system becomes more stable. When $e_{\rm o,eq}>e_{\rm i,eq}$, $f(q)<0$, indicating that the system is dominated by the inner planet and all traps are stable traps.
B.3: Dependence on the mass ratio $q$
We run simulations to testify to our proposed overstability and escape criteria. Figure B.1 shows the numerical results of $\tau_{\rm a}/\tau_{\rm e}$ as a function of $q$. We fix $m_{\rm i}{=}10\ M_{\oplus}$ and set $\tau_{\rm a}{=}3\ (\tau_{\rm a})_{\rm slow,mig}$ to fulfill the trapping criteria. The solid lines denote the analytical criteria, while the colored dots represent the simulation results.
The numerical results agree pretty well with the analytical expectation at $q{\leq}1$. The simulations depict a transition from stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes as $\tau_{\rm a}/\tau_{\rm e}$ decreases, and our analytical criteria separate these three regimes. In contrast, when $q\geq 1$, there are still two points in Figure B.1 where the simulations do not fully match the analysis.
First, the stability/overstability and escape criteria exhibit the opposite dependence $q$, with the two lines intersecting at $q{\sim}2$. For $q{\geq}2$, one would expect that the systems reside region below the overstability criterion (yellow line) correspond to escape. However, simulations show that all systems in this region are trapped overstably (no escape occurs). We attribute this to the fact that when $q{\geq}2$, the angular momentum deficit (AMD) (Laskar 1997) of the inner planet is larger than that of the outer planet, limiting the amplitude of oscillations in overstability (see Appendix B.4 for detailed derivation). As a result, the perturbations cannot grow sufficiently to cause escape from resonance. The dashed purple line in Figure B.1 represents the AMD criterion when the second-order terms of the eccentricities are considered. This new criterion is in good agreement with our numerical simulations.
Second, our overstable criterion suggests $\tau_{\rm a}/\tau_{\rm e}{<}0$ at $q{\geq}3$, indicating that all traps should be stable in this regime. However, in Figure B.1, it is evident that near $q\sim 4$ and $\tau_{a}/\tau_{e}\sim 10^{2}$, the planets remain in an overstability state. This is because we only considered the first-order term of the disturbing function. Second-order terms are neglected here, weakening the inner planet effect on the outer planet and leading to an overstable trap.
Figure B.1: Simulation results of $2:1$ MMR stability in $q-\tau_{a}/\tau_{e}$ parameter space with the mass of the inner planet $m_{\rm i}=10\ \times M_{\oplus}$. Green points indicate stable resonance capture, yellow points indicate overstable resonance capture, and purple points indicate the planets eventually escape from the resonance. The lines represents our analytical criteria. It can be observed that our criteria overall align well with the numerical results.
B.4: Second-Order AMD Criterion
We expand the disturbing functions to the second order in $e_{\rm i}$ and $e_{\rm o}$, which gives
$R_{\rm D}^{\rm sec}$ $=$ $K_{1}(e_{\rm i}^{2}+e_{\rm o}^{2})+K_{2}e_{\rm i}e_{\rm o}\cos{(\varpi_{\rm i}-\varpi_{\rm o})},$ (B8)
$R_{\rm D}^{\rm res}$ $=$ $e_{\rm i}f_{\rm d,i}\cos{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}\cos{\varphi_{\rm o}}$ (B9)
$+[e_{\rm i}^{2}f_{3}\cos{\varphi_{3}}+e_{\rm i}e_{\rm o}f_{4}\cos{\varphi_{4}}+e_{\rm o}^{2}f_{5}\cos{\varphi_{5}}]\delta_{\rm j,2},$
$R_{\rm E}$ $=$ $-2e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2},$ (B10)
$R_{\rm I}$ $=$ $-\frac{1}{2}e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2},$ (B11)
where $\varphi_{3}{=}4\lambda_{\rm o}-2\lambda_{\rm i}-2\varpi_{\rm i}$, $\varphi_{4}{=}4\lambda_{\rm o}-2\lambda_{\rm i}-\varpi_{\rm i}-\varpi_{\rm o}$, $\varphi_{5}{=}4\lambda_{\rm o}-2\lambda_{\rm i}-2\varpi_{\rm o}$. Detailed expressions of $K_{1},K_{2},f_{3},f_{4}$ and $f_{5}$ can be found in Tables 8.1 and 8.4 of Murray & Dermott (1999). Lagrange’s equations can be written as
$\dot{\varpi_{\rm i}}$ $=$ $\frac{\mu_{\rm o}n_{\rm i}\alpha}{e_{\rm i}}[f_{\rm d,i}\cos{\varphi_{\rm i}}+(2e_{\rm i}f_{3}\cos{\varphi_{3}}+e_{\rm o}f_{4}\cos{\varphi_{4}})\delta_{\rm j,2}$ (B12)
$+2K_{1}e_{\rm i}+K_{2}e_{\rm o}\cos{(\varpi_{\rm i}-\varpi_{\rm o})}],$
$\dot{\varpi_{\rm o}}$ $=$ $\frac{\mu_{\rm i}n_{\rm o}}{e_{\rm o}}[f_{\rm d,o}^{{}^{\prime}}\cos{\varphi_{\rm o}}+(2e_{\rm o}f_{5}\cos{\varphi_{5}}+e_{\rm i}f_{4}\cos{\varphi_{4}})\delta_{\rm j,2}$ (B13)
$+2K_{1}e_{\rm o}+K_{2}e_{\rm i}\cos{(\varpi_{\rm i}-\varpi_{\rm o})}].$
The angular momentum deficit (AMD) (Laskar 1997) is given by
$AMD\sim\mu\sqrt{a}(1-\sqrt{1-e^{2}})\sim\mu\sqrt{a}e^{2}.$ (B14)
When the AMD of the inner planet exceeds that of the outer planet, the condition
$e_{\rm o}^{2}/e_{\rm i}^{2}>q\sqrt{\alpha},$ (B15)
is satisfied. Under this condition, the amplitude of oscillations following overstability is limited, and therefore the perturbations cannot grow sufficiently to cause escape. By substituting Eqs. (B12) and (B13) to eliminate either $e_{\rm i}$ and $e_{\rm o}$ under the condition $\dot{\varpi_{\rm i}}=\dot{\varpi_{\rm o}}$, we obtain the relationship between $e_{\rm i,o}$ and $q$. Assuming $e_{\rm i}$ takes the form: $e_{\rm i}^{2}={\tau_{e}}/{2j\tau_{a}}$, and substituting this into the derived form, we obtain our AMD criterion as
$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm AMD}=\frac{2}{j}\left[\frac{(f_{3}-K_{1})+q^{{}^{\prime}2}(f_{5}+K_{1})}{q^{{}^{\prime}}f_{\rm d,o}^{{}^{\prime}}+f_{\rm d,i}}\right]^{2},$ (B16)
where $q^{{}^{\prime}2}\equiv q\sqrt{\alpha}$. The AMD criterion aligns well with the numerical results (Figure B.1).

### v2

@@
 Resonance Capture and Stability Analysis for Planet Pairs under Type I Disk Migration
-Linghong Lin 11    Beibei Liu 11⋆⋆    Zekai Zheng 1122
-We present a theoretical model to investigate a two-planet pair that undergoes convergent type I migration in a gaseous protoplanetary disk and traps into the first-order mean motion resonance.
-Our study identifies the conditions for resonant capture and explores the subsequent dynamical stability of the system.
-We derive the analytical criteria for planets with an arbitrary mass ratio and validate through numerical N-body simulations.
-Slow migration and weak eccentricity damping are required for resonant capture, the latter of which has not received enough attention in the literature.
-Once capture into resonance, their following stability can be classified into three regimes: stable trap, overstable trap and escape.
-Notably, resonant capture remains stable when the inner planet significantly outweighs the outer one.
-However, the subsequent evolution can be diverse when the mass of the inner planet is lower or comparable to that of the outer planet.
-Stability weakens as the relative strength between migration and eccentricity damping increases.
-The key results can be comprehensively demonstrated in a $\tau_{a}$-$\tau_{a}/\tau_{\rm e}$ plot, where $\tau_{a}$ and $\tau_{\rm e}$ are orbital decay and eccentricity damping timescales, respectively.
+Linghong Lin 11    Beibei Liu 1 ⋆1 ⋆    Zekai Zheng 1122
+We present a theoretical framework for investigating a two-planet system undergoing convergent type I migration in a protoplanetary disk.
+Our study identifies the conditions for resonant capture and subsequent dynamical stability.
+By deriving analytical criteria for general $j$:j1j-1 first-order mean-motion resonances (MMRs) applicable to planet pairs with arbitrary mass ratios, we validate these predictions through N-body simulations.
+The key results are demonstrated in $\tau_{\rm m}$-$\tau_{\rm m}/\tau_{e}$ plots, where $\tau_{\rm m}$ and $\tau_{e}$ are the timescales of the angular momentum and eccentricity damping, respectively.
+Specifically, we determine which combinations of orbital damping timescales allow for capture into resonance, showing that too fast migration or too strong eccentricity damping inhibit successful capture.
+After capture, the subsequent evolution can be classified into three regimes: stable trap, overstable trap and escape.
+Importantly, resonant capture always remains stable when the inner planet significantly outweighs the outer one.
+In contrast, when the mass of the inner planet is lower than or comparable to that of the outer planet, the system transitions from the stable to overstable trap, and eventually escapes the resonance, as the relative strength of eccentricity damping to migration ($\tau_{\rm m}/\tau_{e}$) decreases.
 Key Words.:
@@
 1 Introduction
-Low-mass planets embedded in gaseous protoplanetary disks interact with disk gas and undergo type I orbital migration (Kley & Nelson, 2012; Paardekooper et al., 2023).
-Resonance capture can be a natural outcome of the planet pairs featured with convergent migration (Lee & Peale, 2002).
-The key question is to understand the under which conditions this capture can occur.
-The classical pendulum model pointed out that the trapping holds when the time planet migrates across the resonance width is longer than the time resonant angle librates (Murray & Dermott, 1999; Ogihara & Kobayashi, 2013; Batygin, 2015; Liu et al., 2015).
-A new criterion was derived by Huang & Ormel (2023), where they analyzed the resonant equilibrium state under the disk migration effects in a restrict three-body framework.
-Batygin & Petit (2023) provided an equivalent criterion using the Hamiltonian theory and extended it to the context of the unrestricted problem.
-On the other hand, Goldreich & Schlichting (2014) first investigated the long-term stability of the resonant pair after capture in a gaseous disk.
-They found that under certain conditions, the resonant equilibrium state can become overstable, indicating that the planets may temporarily or permanently escape from resonance.
-Deck & Batygin (2015) conducted a similar analysis based on the Hamiltonian theory.
-Xu & Lai (2017) extended the framework to second-order mean motion resonances.
-Their study highlighted that the mass ratio of the planet pair is a crucial parameter in determining the overstability.
+Low-mass planets embedded in protoplanetary disks interact with surrounding disk gas and undergo type I orbital migration (Goldreich & Tremaine, 1980; Ward, 1997; Tanaka et al., 2002; Kley & Nelson, 2012; Baruteau et al., 2014; Paardekooper et al., 2023).
+Resonance capture can be a natural outcome of a planet pair that undergoes convergent migration (Lee & Peale, 2002; Papaloizou & Szuszkiewicz, 2005).
+The key question is to understand under which conditions this capture can occur and its following dynamical evolution.
+The classical pendulum model intuitively shows that the trapping holds when the time planet migrates across the resonance width is longer than the libration period in the resonance (Murray & Dermott, 1999; Ogihara & Kobayashi, 2013; Liu et al., 2015), which is akin to the adiabatic condition described by Batygin (2015).
+The eccentricity damping from the gas disk is neglected in these studies.
+Later, Kajtazi et al.
+(2023) demonstrated through simulations that eccentricity damping qualitatively alters the capture conditions.
+Incorporating eccentricity damping, Huang & Ormel (2023) derived a new criterion by analyzing the resonant equilibrium state under disk-driven migration within a restricted three-body framework.
+Meanwhile, Batygin & Petit (2023) derived an equivalent criterion using Hamiltonian theory and extended the analysis to the two-planet pairs with arbitrary mass ratios.
+On the other hand, Goldreich & Schlichting (2014) first investigated the long-term stability of the resonant planet pair after capture in a gaseous disk.
+Their study revealed that under certain conditions, the resonant equilibrium state can become overstable, where the planets may eventually escape from resonance.
+Building on this, Deck & Batygin (2015) conducted a similar analysis using Hamiltonian theory and extended the work to the unrestricted three-body system.
+Xu & Lai (2017) further expanded the framework to second-order mean motion resonances.
+These studies highlighted that the mass ratio of the planet pair is a crucial parameter in determining whether the system is overstable.
 Nesvorný et al.
@@
 (2024) verified this overstable resonant trap using hydrodynamical simulations.
-They found that the outcome is related to the disk properties, such as disk mass and viscosity.
-In this paper, we develop a unified framework for a two-planet pair that undergoes convergent type I migration and gets trapped at the first-order mean motion resonance ($j$:$j-1$ MMR).
-We investigate the conditions under which planets can be captured in resonances and examine the following stability properties.
-The key analytical criteria are derived for planets with arbitrary mass ratios and the outcomes are summarized in a $\tau_{\rm a}{-}\tau_{\rm a}/\tau_{\rm e}$ diagram.
+They found that the outcome is related to disk properties, such as the gas mass and turbulent viscosity.
+In this paper, we develop a unified framework for a two-planet pair that undergoes convergent type I migration and gets trapped at the first-order mean motion resonance ($j$:j1j-1 MMR).
+We investigate the conditions under which planets can be captured in resonances and examine the subsequent stability properties.
+The key analytical criteria are derived for planets with arbitrary mass ratios, and the outcomes are summarized in a $\tau_{\rm m}{-}\tau_{\rm m}/\tau_{e}$ parameter space.
 The paper is organized as follows.
-We present an analytical model accounting for the dynamics of planets at resonances in Sec.
-2.
-In Sec.
-3 we perform N-body simulations and compare the results with the analytical predictions.
-In Sec.
-4 we discuss several implications as well as potential caveats in our model.
-Finally, we summarize the key results in Sec.
-5.
-2 Dynamics of resonance trapping
-In this section, we discuss the dynamical evolution of a two-planet pair in a gaseous protoplanetary disk after their convergent type-I migration.
+In Section 2, we present an analytical model that describes the dynamics of planets in resonant configurations.
+In Section 3 we perform N-body simulations and compare the results with analytical predictions.
+Section 4 outlines potential caveats, and Section 5 summarizes our key results.
+2 Dynamics of resonance capture and stability
+This section discusses the dynamical evolution of a two-planet pair in a gaseous disk through the convergent type-I migration.
 We consider two planets of masses $m_{\rm i}$ and $m_{\rm o}$ orbiting a central star of $M_{\star}$ on coplanar orbits.
-Their reduced masses and the mass ratio between the inner and outer ones are $\mu_{\rm i,o}{=}m_{\rm i,o}/M_{\rm\star}$ and $q{=}m_{\rm i}/m_{\rm o}$, respectively.
-For notations the subscripts ‘i’ and ‘o’ refer to the corresponding quantities of inner and outer planets.
-We present the analytical criteria of resonance trapping and stability in Sec. 2.1 and Sec. 2.2, respectively. A unified picture is summarized in Sec. 2.3.
+Their normalized masses and the mass ratio between the inner and outer ones are $\mu_{\rm i,o}{=}m_{\rm i,o}/M_{\rm\star}$ and $q{=}m_{\rm i}/m_{\rm o}$, respectively.
+For notation, the subscripts ‘i’ and ‘o’ refer to the corresponding quantities of the inner and outer planets.
+We present the analytical criteria for resonance trapping and stability in Section 2.1 and Section 2.2, respectively.
+A unified picture is summarized in Section 2.3.
+Resonance crossing in the context of divergent migration is also discussed in Section 2.4.
 2.1 Resonance trapping criteria
-The dissipative effect of planet-disk interaction can be parameterized by (Teyssandier & Terquem, 2014)
-$\frac{1}{a}\frac{da}{dt}$ $=$ $-\frac{1}{\tau_{a}}-\frac{2e^{2}}{\tau_{e}},$ (1a)
-$\frac{1}{e}\frac{de}{dt}$ $=$ $-\frac{1}{\tau_{e}},$ (1b)
-where $\tau_{a}$ and $\tau_{e}$ are the orbital decay and eccentricity damping timescales, respectively.
-The orbital elements $a$, $n$, $e$, $\lambda$, $\varpi$, denote the semimajor axis, mean motion, eccentricity, mean longitude and longitude of pericenter of the planet.
+The dissipative effect of the planet-disk interaction can be parameterized by (Teyssandier & Terquem, 2014; Ataiee & Kley, 2021; Pichierri et al., 2023)
+$\frac{1}{L}\frac{dL}{dt}$ $=$ $-\frac{1}{\tau_{\rm m}},$ (1a)
+$\frac{1}{a}\frac{da}{dt}$ $=$ $-\frac{1}{\tau_{\rm a}}=-\frac{2}{\tau_{\rm m}}+\frac{2e^{2}}{1-e^{2}}\frac{1}{\tau_{e}},$ (1b)
+$\frac{1}{e}\frac{de}{dt}$ $=$ $-\frac{1}{\tau_{e}},$ (1c)
+where the orbital elements $L$, $a$, $n$, $e$, $\lambda$, $\varpi$, denote the angular momentum, semimajor axis, mean motion, eccentricity, mean longitude and longitude of pericenter of the planet, $\tau_{\rm m}$, $\tau_{\rm a}$ and $\tau_{\rm e}$ represent the damping timescales for orbital angular momentum, semi-major axis and eccentricity, respectively.
+It is worth noting that in the limit of small eccentricity, the angular momentum and orbital decay timescales can be linked through $\tau_{\rm m}{\approx}2\tau_{\rm a}$.
 Figure 1: Resonance crossing of a two-planet system with convergent migration.
-The plots provide the time evolution of the outer-to-inner planet period ratio (panel a), the two planets’ eccentricities (panel b), resonant angle (panel c) and phase curve (panel d).
-The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, the planet masses are $m_{\rm i}=1\ M_{\oplus}$, $m_{\rm o}=10\ M_{\oplus}$, and the migration parameters are $\tau_{a}=1.1\times 10^{5}yr$ and $\tau_{a}/\tau_{e}=1.5\times 10^{3}$.
-The planets directly cross the $2$:$1$ resonance at $t{\simeq}6\times 10^{3}$ yr and their eccentricities simultaneously excite to $10^{-2}$ and rapidly decline afterward.
-The resonant angle of the inner planet jumps from the initial values to $\pi$ during the crossing and oscillates over $2\pi$ afterwards.
-In phase space, the inner planet is not captured by the non-zero resonant fixed point and ultimately evolves towards the origin.
-For the $j$:$j-1$ order mean motion resonances (MMRs), the resonant angles and the time-averaged disturbing functions for the inner and outer planets are given by (Murray & Dermott, 1999) $\varphi_{\rm i}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm i}$, $\varphi_{\rm o}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm o}$, and $R_{\rm i}{=}Gm_{\rm o}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+\alpha R_{\rm E})$, $R_{\rm o}{=}Gm_{\rm i}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+R_{\rm I}/\alpha^{2})$, where $\alpha{=}a_{\rm i}/a_{\rm o}$ is the two planets’ semi-major axis ratio, $R_{\rm D}^{\rm sec}$ and $R_{\rm D}^{\rm res}$ are the secular and resonant terms, $R_{\rm E}$ and $R_{\rm I}$ are the indirect part due to an external and internal perturber.
-We assume small eccentricities of the planets, and expand the disturbing functions to the first order in $e_{\rm i}$ and $e_{\rm o}$, which gives $R_{\rm D}^{\rm sec}{=}0$, $R_{\rm D}^{\rm res}{=}e_{\rm i}f_{\rm d,i}\cos{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}\cos{\varphi_{\rm o}}$,$R_{\rm E}{=}-2e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2}$, and $R_{\rm I}{=}-\frac{1}{2}e_{\rm o}\cos{\varphi_{\rm o}}\delta_{\rm j,2}$, where $\delta_{\rm j,2}$ is the usual Kronecker symbol.
+The plot shows the time evolution of (a) the outer-to-inner planet period ratio, (b) the planet eccentricities, (c) the resonant angle, and (d) the phase curve.
+The planet and disk parameters are $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $\tau_{\rm m}{=}2.2\times 10^{5}$ yr and $\tau_{\rm m}/\tau_{e}{=}3\times 10^{3}$.
+The planets directly cross the 22:11 resonance at $t{\simeq}6\times 10^{3}$ yr, while the inner planet’s eccentricity briefly excites to $0.01$ before rapid decay.
+During the crossing, the resonant angle of the inner planet $\varphi_{\rm i}$ jumps from 0 to $\pi$ and subsequently oscillates with growing amplitude22footnotemark: 2.
+In phase space, the planet’s trajectory is not captured by the resonant equilibrium.
+Instead, it circulates and eventually evolves toward the origin, corresponding to a low-eccentricity, non-resonant state.
+See https://github.com/llh-astro/resonance/raw/main/gif/no_trap.gif
+For a $j$:j1j-1 MMR, the resonant angles and the time-averaged disturbing functions for the inner and outer planets are given by (Murray & Dermott, 1999) $\varphi_{\rm i}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm i}$, $\varphi_{\rm o}{=}j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm o}$, and $R_{\rm i}{=}Gm_{\rm o}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+\alpha R_{\rm E})$, $R_{\rm o}{=}Gm_{\rm i}/a_{\rm o}(R_{\rm D}^{\rm sec}+R_{\rm D}^{\rm res}+R_{\rm I}/\alpha^{2})$, where $\alpha{=}a_{\rm i}/a_{\rm o}$ is the two planets’ semi-major axis ratio, $R_{\rm D}^{\rm sec}$ and $R_{\rm D}^{\rm res}$ are the secular and resonant terms, $R_{\rm E}$ and $R_{\rm I}$ are the indirect part due to an external and internal perturber.
+We assume small eccentricities of the planets, and expand the disturbing functions to the first order in $e_{\rm i}$ and $e_{\rm o}$, which gives $R_{\rm D}^{\rm sec}{=}0$, $R_{\rm D}^{\rm res}{=}e_{\rm i}f_{\rm d,i}\cos{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}\cos{\varphi_{\rm o}}$$R_{\rm E}{=}-2\delta_{\rm j,2}e_{\rm o}\cos{\varphi_{\rm o}}$, and $R_{\rm I}{=}-\frac{1}{2}\delta_{\rm j,2}e_{\rm o}\cos{\varphi_{\rm o}}$, where $\delta_{\rm j,2}$ is the usual Kronecker symbol.
 Detailed expressions of $f_{\rm d,i}$ and $f_{\rm d,o}$ can be found in Table 8.1 of Murray & Dermott (1999).
 Taking orbital decay and eccentricity damping into account, Lagrange’s equations can be written as
-$\dot{n_{\rm i}}$ $=-3(j-1)\mu_{\rm o}n_{\rm i}^{2}\alpha(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm i}}{2\tau_{a,\rm i}}+\frac{3n_{\rm i}e_{\rm i}^{2}}{\tau_{e,\rm i}},$ (2a)
-$\dot{n_{\rm o}}$ $=3j\mu_{\rm i}n_{\rm o}^{2}(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm o}}{2\tau_{a,\rm o}}+\frac{3n_{\rm o}e_{\rm o}^{2}}{\tau_{e,\rm o}},$ (2b)
+$\dot{n_{\rm i}}$ $=-3(j-1)\mu_{\rm o}n_{\rm i}^{2}\alpha(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm i}}{\tau_{\rm m\rm,i}}+\frac{3n_{\rm i}e_{\rm i}^{2}}{\tau_{e,\rm i}},$ (2a)
+$\dot{n_{\rm o}}$ $=3j\mu_{\rm i}n_{\rm o}^{2}(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})+\frac{3n_{\rm o}}{\tau_{\rm m\rm,o}}+\frac{3n_{\rm o}e_{\rm o}^{2}}{\tau_{e,\rm o}},$ (2b)
 $\dot{e_{\rm i}}$ $=-\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}\sin{\varphi_{\rm i}}-\frac{e_{\rm i}}{\tau_{e,\rm i}},$ (2c)
@@
 $\dot{\varpi_{\rm o}}$ $=\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}}\frac{\cos{\varphi_{\rm o}}}{e_{\rm o}},$ (2f)
-where $f_{\rm d,o}^{{}^{\prime}}{=}f_{\rm d,o}-2\alpha\delta_{j,2}$. For instance, $f_{\rm d,i}{=}{-}1.19$, $f_{\rm d,o}^{{}^{\prime}}{=}0.43$ at the $2$:$1$ MMR.
-When the planets are in MMRs, their orbital parameters reach the equilibrium state such that $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$, $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$.
-By substituting $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$ into Laplace equation, Eqs.
+where $f_{\rm d,o}^{{}^{\prime}}{=}f_{\rm d,o}-2\alpha\delta_{j,2}$. For instance, $f_{\rm d,i}{=}{-}1.19$, $f_{\rm d,o}^{{}^{\prime}}{=}0.43$ at the 22:11 MMR.
+The orbital parameters of the planets reach the equilibrium state at resonances such that $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$, $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$.
+By substituting $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$ into Lagrange’s equation, Eqs.
 (2a) and (2b) can be combined as
 $-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})$ (3)
-$=\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
-where we assume that $\tau_{a}^{-1}{=}\tau_{a,\rm o}^{-1}-\tau_{a,\rm i}^{-1}$ is the relative orbital decay rate with respect to convergent migration, and $\tau_{e}{=}\tau_{e,\rm i}{=}\tau_{e,\rm o}$ represents the eccentricity damping time-scale.
+$=\frac{1}{\tau_{\rm m}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
+where we define that $\tau_{\rm m}^{-1}{=}\tau_{\rm m,o}^{-1}-\tau_{\rm m,i}^{-1}$ is the relative angular momentum damping rate for convergent migration, and $\tau_{e}=\tau_{e,\rm i}=\tau_{e,\rm o}$ represents the eccentricity damping timescale.
+In reality, the eccentricity damping rates of the two planets may vary, depending on disk properties and planet masses.
+For simplicity, we assume equal $\tau_{e}$ for both planets in our theoretical derivations and numerical simulations in the main text.
+We note, however, that our theoretical framework can be naturally extended to the case of unequal eccentricity damping.
+The expressions of our criteria for different eccentricity damping timescales are elaborated in Appendix C.
 Substituting $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ into Eqs.
 (2c) and (2d) to eliminate $\sin{\varphi_{\rm i,o}}$ and $\dot{\varpi_{\rm i}}{=}\dot{\varpi_{\rm o}}$ into Eqs.
-(2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$, the equilibrium eccentricities can be derived from Eq.
-(3) such that
-$e_{\rm i,eq}^{2}$ $=$ $\frac{\tau_{e}}{2\tau_{a}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4a)
-$e_{\rm o,eq}^{2}$ $=$ $\frac{\tau_{e}}{2\tau_{a}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4b)
-where $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$.
-Our results are analogous to Terquem & Papaloizou (2019), except for $\tau_{a}$ and $\tau_{e}$ as defined.
-In the limit of the mass of inner planet much lower than the outer one ($q{\ll}1$), the eccentricity of inner planet in Eq.
-(4a) can be simplified as
-$e_{\rm i,eq}=\sqrt{\frac{1}{2j}\frac{\tau_{e}}{\tau_{a}}},$ (5)
-in agreement with Eq.
-(24) in Huang & Ormel (2023).
-We note that due to differing signs of $f_{\rm d,i}$ and $f_{\rm d,o}^{{}^{\prime}}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ requires that $\sin{\varphi_{\rm i}}{>}0$ and $\sin{\varphi_{\rm o}}{<}0$ for $j$:$j-1$ MMRs.
+(2e) and (2f) to eliminate either $e_{\rm i}$ or $e_{\rm o}$.
+Here we assume $\cos{\varphi_{\rm i}}=\cos{\varphi_{\rm o}}=1$.
+Under these conditions, the equilibrium eccentricities can be derived from Eq.
+(3), yielding
+$e_{\rm i,eq}^{2}$ $=$ $\frac{\tau_{e}}{\tau_{\rm m}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4a)
+$e_{\rm o,eq}^{2}$ $=$ $\frac{\tau_{e}}{\tau_{\rm m}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (4b)
+where $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$ (see Lee 2004; Deck & Batygin 2015; Ramos et al. 2017; Terquem & Papaloizou 2019).
+In the limit of the mass of the inner planet much lower than the outer one ($q{\ll}1$), the eccentricity of the inner planet in Eq. (4a) can be simplified as
+$e_{\rm i,eq}=\sqrt{\frac{1}{j}\frac{\tau_{e}}{\tau_{\rm m}}}.$ (5)
+We note that due to differing signs of $f_{\rm d,i}$ and $f_{\rm d,o}^{{}^{\prime}}$, $\dot{e_{\rm i}}{=}\dot{e_{\rm o}}{=}0$ requires that $\sin{\varphi_{\rm i}}{>}0$ and $\sin{\varphi_{\rm o}}{<}0$ for $j$:j1j-1 MMRs.
 Considering $e_{\rm o,eq}/e_{\rm i,eq}{=}{q\sqrt{\alpha}|f_{\rm d,o}^{{}^{\prime}}/f_{\rm d,i}|}$, we can derive $\sin{\varphi_{\rm i}}{=}-\sin{\varphi_{\rm o}}$, which implies $\varphi_{\rm i}-\varphi_{\rm o}{=}\pm\pi$.
-This means that the conjunctions between planet pair in a mean motion resonance always occur at either pericentre or apocentre.
-Notably, the right-hand side of Eq.
-(3) represents the planet-disk interaction that leads convergent orbital migration between two planets, while the left-hand side of Eq.
-(3) represents their resonant gravitational interaction, keeping them away from getting closer.
+This means that the conjunctions between planet pairs in a mean motion resonance always occur at either pericentre or apocentre.
+In particular, the right-hand side of Eq.
+(3) represents the planet-disk interaction that leads to convergent orbital migration between two planets, while the left-hand side of Eq.
+(3) represents their resonant gravitational interaction.
 When the effect of planet-disk interaction is strong enough to overcome the resonant interaction, the planet pair cannot stay in resonance.
-Strong planet-disk interaction can be further classified into two regimes: fast migration and strong eccentricity damping.
-In the first regime the exchange rate of semimajor axis $\dot{a}/a$ in Eq.
-(3) is driven by the a-decay term ($\tau_{a}^{-1}{\gg}e^{2}\tau_{e}^{-1}$).
-In the second regime, however, the contribution of the e-damping term becomes comparable to that of the a-decay term.
-We will derive the resonance trapping conditions in the following subsections.
-Figure 2: Resonance crossing of a two-planet system with convergent migration.
-The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, and the migration parameters are $\tau_{a}=10^{5}yr$ and $\tau_{a}/\tau_{e}=6\times 10^{2}$.
-The planets are trapped in the $2:1$ resonance at $t{\simeq}10^{4}$ yr and their eccentricities simultaneously excite to $2\times 10^{-2}$, which remain stable thereafter.
-The resonant angle of the inner planet is excited when the period ratio evolves to 2:1, subsequently oscillating and gradually converging to the resonant equilibrium value, which it maintains.
-In phase space, the inner planet is captured by the non-zero resonant fixed point and subsequently evolves in conjunction with this fixed point, reflecting the static equilibrium state of the resonance..
+The strong planet-disk interaction can be classified into two regimes: fast migration and strong eccentricity damping.
+In the first regime, $\dot{a}/a$ in Eq.
+(3) is driven by the angular momentum damping term ($\tau_{\rm m}^{-1}{\gg}e^{2}\tau_{e}^{-1}$).
+In the second regime, however, the contribution of the e-damping term becomes comparable to that of the angular momentum damping term.
+We will derive the resonance trapping conditions from these two regimes in the following subsections.
+Figure 2: Stable resonant capture of a two-planet system with convergent migration.
+The planet and disk parameters are $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $\tau_{\rm m}{=}2\ \times 10^{5}$ yr and $\tau_{\rm m}/\tau_{e}{=}1200$.
+The planets get trapped in a 22:11 MMR at $t{\simeq}10^{4}$ yr exciting the inner planet’s eccentricity to $0.02$.
+The resonant angle $\varphi_{\rm i}$ shifts from 0 to $2\pi/3$ as the system transitions into the resonant state.
+It gradually decreases as convergent migration strengthens the resonant interaction, causing the libration center to drift slowly.
+In phase space, the trajectory converges toward the resonant equilibrium and co-evolves with it.
+We refer to this resonant capture at a stationary equilibrium as a stable trap.
+See https://github.com/llh-astro/resonance/raw/main/gif/stable_trap.gif
 2.1.1 Slow migration
-When the planets undergo fast convergent migration, they directly bypass the resonance before the eccentricities reach to the equilibrium values.
-In this case $\dot{a}/a$ is driven by $\tau_{a}$ but not $\tau_{e}$, and Eq.
+Relatively slow migration is a necessary resonant capture condition.
+When the planets undergo fast convergent migration, they directly bypass the resonance before the eccentricities reach their equilibrium values.
+In this case $\dot{a}/a$ is driven by $\tau_{\rm m}$ but not $\tau_{e}$, and Eq.
 (3) can be rewritten as
-$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{2\tau_{a}},$ (6)
+$(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{\tau_{\rm m}},$ (6)
 where the e-damping term is ignored.
-Here the eccentricities are purely excited by the resonant term, which can be deduced from the classical pendulum model (see appendix Sec.
-Appendix A: The derivation of the excited eccentricity in the classical pendulum model for derivation)
+We adopt $\sin{\varphi_{\rm i}}{=}-\sin{\varphi_{\rm o}}{=}1$, so the left-hand side of Eq.
+6 represents the maximum strength of the resonant interaction.
+From the classical pendulum model (see Appendix A for the derivation), the maximum value of $e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}$ can be obtained as
 $e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}=-\frac{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}.$ (7)
-Combing Eqs. (6) and (7), the resonance trapping criterion of slow migration can be obtained as
-$\tau_{a}>\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}.$ (8)
+If the migration rate is too rapid, the migration term in Eq.
+6 always exceeds the resonant term, and therefore the equilibrium state cannot be reached.
+Combing Eqs.
+(6) and (7), the resonance trapping criterion of slow migration can be written as
+$\tau_{\rm m}>\frac{1}{[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}.$ (8)
 We further simplify the above criterion as
-$\begin{cases}{\displaystyle\tau_{a}>\frac{1}{2n_{\rm i}(3(j-1))^{1/3}}\frac{1}{(\mu_{\rm o}\alpha f_{\rm d,i})^{4/3}}\ \sim\mu_{\rm o}^{-4/3}}\ \ \hfill\mbox{ at $q\ll 1$},\\ {\displaystyle\tau_{a}>\frac{1}{2n_{\rm o}(3j)^{1/3}}\frac{1}{(\mu_{\rm i}f_{\rm d,o}^{{}^{\prime}})^{4/3}}\ \sim\mu_{\rm i}^{-4/3}}\ \ \ \hfill\mbox{at $q\gg 1$}.\\ \end{cases}$ (9)
+[…]
 This indicates that the threshold migration rate increases with the planet mass.
-Intuitively, the planet resonant interaction becomes stronger as their masses increase.
+Intuitively, the planet’s resonant interaction becomes stronger as their masses increase.
 Therefore, the planets require faster migration in order to break the resonance.
 2.1.2 Weak eccentricity damping
-When the planets migrate relatively slow, they have time to reach the equilibrium state.
+When the planets migrate relatively slowly towards each other, they have time to reach the equilibrium state.
 Substituting the equilibrium eccentricity into the right side of Eq.
-(3), it turns out that the magnitude of e-damping term is comparable to that of the a-decay term.
+(3), it turns out that the magnitude of the e-damping term is comparable to that of the angular momentum damping term.
 In this circumstance, the e-damping term becomes a key factor for resonance trapping.
-Considering the limit of $\sin{\varphi_{\rm i}}{=}1$ and $\sin{\varphi_{\rm o}}{=}-1$, Eq. (3) can be rewritten as
-$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$ (10)
-where the left-hand side of the planetary interaction maintains a finite magnitude at resonances (${\propto}\mu ne$).
-If eccentricity damping is very strong, the orbital circularization leads large $\dot{a}/a$,, and the resonance can no longer hold.
+By substituting $e_{\rm i,eq}$ and $e_{\rm o,eq}$ into Eqs.(2c) and (2d), $\sin{\varphi_{\rm i,o}}$ can be calculated.
+Naturally, the absolute values of $\sin{\varphi_{\rm i,o}}$ cannot exceed $1$.
+Therefore, by considering the limit of $\sin{\varphi_{\rm i}}{=}1$ and $\sin{\varphi_{\rm o}}{=}-1$, Eq.(3) can be rewritten as
+$(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})=\frac{1}{\tau_{\rm m}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}}.$ (10)
+If eccentricity damping is very strong, the orbital circularization leads to large $\dot{a}/a$, and the resonance can no longer hold.
 One example of resonance crossing is shown in Figure 2.
@@
 (4a) and (4b), we obtain the resonance trapping criterion in the weak e-damping regime as
-$\tau_{e}\tau_{a}>\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}.$ (11)
-In the limit of $q{\ll}1$, the trapping criterion can be simplified as
-$\tau_{e}\tau_{a}>\frac{1}{2j(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i})^{2}},$ (12)
-which is in consistent with Eq. (25) of Huang & Ormel (2023).
-Altogether, the resonance trapping requires both relatively weak eccentricity damping and slow migration.
+$\tau_{e}\tau_{\rm m}>\frac{1}{(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}.$ (11)
+The expression agrees with the one obtained by Batygin & Petit (2023), who used a different yet equivalent analytical method. In the limit of $q{\ll}1$, the trapping criterion can be simplified as
+$\tau_{e}\tau_{\rm m}>\frac{1}{j(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i})^{2}},$ (12)
+consistent with Eq. (25) of Huang & Ormel (2023).
+Altogether, resonance trapping requires both relatively weak eccentricity damping and slow migration.
 Taking Eqs.
-(11) and (8) together, we formulate the necessary condition for $\tau_{a}$ and $\tau_{e}$ as
-$\begin{cases}{\displaystyle(\tau_{e}\tau_{a})_{\rm weak,e}=\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vspace{0.1cm}\\ {\displaystyle(\tau_{a})_{\rm slow,mig}=\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}].\end{cases}$ (13)
-Figure 3: Resonance crossing of a two-planet system with convergent migration.
-The initial period ratio of the two planets is $2.2$, just outside of the 2:1 MMR, and the migration parameters are $\tau_{a}=4\times 10^{5}yr$ and $\tau_{a}/\tau_{e}=5\times 10^{2}$.
-The planets are trapped in the $2:1$ resonance at $t{\simeq}3\times 10^{4}$ yr and their eccentricities first excite to $2\times 10^{-2}$, and starts to exhibit significant oscillations around its equilibrium value.
-The resonant angle of the inner planet remains at the equilibrium value and begins to oscillate after some time.
-In phase space, the inner planet initially evolves to a non-zero eccentricity resonant fixed point and subsequently starts oscillating around it.
-Over time, the system settles into a limit cycle around the fixed point, reflecting the dynamic equilibrium state of the resonance.
+(11) and (8) together, we formulate the threshold condition for $\tau_{\rm m}$ and $\tau_{e}$ as
+$\begin{cases}{\displaystyle\tau_{e}\tau_{\rm m}>\frac{1}{(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vskip 2.84544pt\\ {\displaystyle\tau_{\rm m}>\frac{1}{[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}].\end{cases}$ (13)
+Figure 3: Overstable resonance capture of a two-planet system with convergent migration.
+The planet and disk parameters are $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $\tau_{\rm m}{=}8\times 10^{5}$ yr and $\tau_{\rm m}/\tau_{e}{=}10^{3}$.
+By $t{\simeq}3\times 10^{4}$ yr, the planets become trapped in a 22:11 MMR.
+The inner planet’s eccentricity rapidly excites to $0.02$, followed by oscillations with significant amplitude, while the resonant angle exhibits similar behavior.
+In phase space, the trajectory initially evolves toward a resonant equilibrium point (black dot) before stabilizing into a limit cycle around this point.
+We term this phenomenon-where the system remains trapped by librating around an equilibrium with the finite amplitude-an overstable trap.
+See https://github.com/llh-astro/resonance/raw/main/gif/overstable_trap.gif
 2.2 Post-evolution after resonant trapping
-Here we assume that the planet pairs have satisfied the resonance trapping conditions in Sec.
-LABEL:sec:riteria and study their following dynamical evolution.
+Here we assume that the planet pairs have satisfied the resonance trapping conditions in Section 2.1 and study their following dynamical evolution.
 The goal is to explore under which configuration the resonance state is temporary or long-term stable.
@@
 We first focus on the situation that the mass of the inner planet is much lower than the outer one ($m_{\rm i}{\ll}m_{\rm o}$, $q{\ll}1$).
-In the absence of planet-disk interactions, the system’s Hamiltonian $H$ near the $j$:$j-1$ MMR can be expressed as
-$-H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (14)
+In the absence of planet-disk interactions, the system’s Hamiltonian $H$ near the $j$:j1j-1 MMR can be expressed as
+$H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (14)
 where
 $\varphi$ $=$ $j\lambda_{\rm o}+(1-j)\lambda_{\rm i}-\varpi_{\rm i},$ (15a)
-$\Phi$ $=$ $[\frac{3(j-1)^{2}}{f_{\rm d,i}\mu_{\rm o}\alpha_{\rm 0}}]^{2/3}\frac{1-\sqrt{1-e_{\rm i}^{2}}}{2}\sim\mu_{\rm o}^{2/3}e_{\rm i}^{2},$ (15b)
-$\eta$ $=$ $[\frac{1}{\sqrt{3}(j-1)f_{\rm d,i}\mu_{\rm o}}]^{2/3}[(j-1)\alpha_{\rm 0}^{-2/3}-j\alpha_{\rm 0}^{5/6}].$ (15c)
-The first two parameters are a pair of conjugate coordinate and momentum, where $\varphi$ is the resonant angle and $\Phi$ represents the angular momentum deficit (AMD).
-The third parameter, $\eta$ is the dimensionless that quantifies the deviation level from the exact resonance.
+$\Phi$ $=$ $\left[\frac{3(j-1)^{2}}{f_{\rm d,i}\mu_{\rm o}\alpha_{\rm 0}}\right]^{2/3}\frac{1-\sqrt{1-e_{\rm i}^{2}}}{2}\sim\mu_{\rm o}^{2/3}e_{\rm i}^{2},$ (15b)
+$\eta$ $=$ $\left[\frac{1}{\sqrt{3}(j-1)f_{\rm d,i}\mu_{\rm o}}\right]^{2/3}[(j-1)\alpha_{\rm 0}^{-2/3}-j\alpha_{\rm 0}^{5/6}].$ (15c)
+The first two parameters are a pair of conjugate coordinates and momentum, where $\varphi$ is the resonant angle and $\Phi$ represents the angular momentum deficit (AMD).
+The third parameter, $\eta$, is dimensionless and quantifies the deviation level from the exact resonance.
 The orbital period ratio of the two planets is the exact integer ratio at $\eta{=}0$.
-The initial semi-major axis ratio of the two planet $\alpha_{\rm 0}\equiv\alpha[1+(j-1)e_{\rm i}^{2}]$ is the conjugate momentum of a fast angle, which can be derived from Eq.8-92 of Murray & Dermott (1999).
+The initial semi-major axis ratio of the two planets $\alpha_{\rm 0}\equiv\alpha[1+(j-1)e_{\rm i}^{2}]$ is the conjugate momentum of a fast angle, which can be derived from Eq.8-92 of Murray & Dermott (1999).
 We note that in the absence of planet-disk interactions, the parameters $\eta$ and $\alpha_{0}$ remain constant.
-As we will demonstrate later, the effects of planet-disk interaction ($\tau_{\rm a},\tau_{\rm e}$) determines the stability of the planet pair in resonance.
 2.2.1 stability/overstability criterion
 When accounting for the planet-disk interaction, the equations of motion of the system can be obtained from
-$\dot{\Phi}{=}-\frac{\partial H}{\partial\varphi}+\frac{\partial\Phi}{\partial e}\frac{de}{dt},\ \ \ \dot{\varphi}{=}\frac{\partial H}{\partial\Phi}.$ (16)
-By incorporating the orbital decay and eccentricity damping terms in Eq.(1), Eq. (16) can be derived as
+$\dot{\Phi}{=}-\frac{\partial H}{\partial\varphi}+\frac{\partial\Phi}{\partial e}\frac{de}{dt}+\frac{\partial\Phi}{\partial a}\frac{da}{dt},\ \ \ \dot{\varphi}{=}\frac{\partial H}{\partial\Phi}.$ (16)
+Since $\Phi$ is independent of $a$ ($\alpha_{0}$ is a constant), the term involving $\partial\Phi/\partial a=0$ can be neglected.
+By incorporating the angular momentum damping and eccentricity damping terms in Eq.(1), Eq.
+(16) can be derived as
 $\dot{\Phi}$ $=$ $\Phi^{1/2}\sin{\varphi}-\frac{2}{\tau_{e}}\Phi,$ (17a)
 $\dot{\varphi}$ $=$ $-\eta-2\Phi+\frac{1}{2\sqrt{\Phi}}\cos{\varphi},$ (17b)
-$\dot{\eta}$ $=$ $\frac{\partial\eta}{\partial\alpha_{0}}\frac{d\alpha_{0}}{dt}{=}-\frac{\beta}{\tau_{a}}+\frac{2(p+2j-2)\Phi}{(j-1)\tau_{e}},$ (17c)
+$\dot{\eta}$ $=$ $\frac{\partial\eta}{\partial\alpha_{0}}\frac{d\alpha_{0}}{dt}{=}-\frac{2\beta}{\tau_{\rm m}}+\frac{4j\Phi}{(j-1)\tau_{e}},$ (17c)
 where
@@
 The system reaches an equilibrium state at resonances when $\dot{\Phi}{=}0$, $\dot{\varphi}{=}0$, $\dot{\eta}{=}0$. Thus, the resonant parameters, to the lowest order, can be written as
-$\Phi_{\rm eq}$ $=$ $\frac{(j-1)\beta\tau_{e}}{4j\tau_{a}},$ (19a)
+$\Phi_{\rm eq}$ $=$ $\frac{(j-1)\beta\tau_{e}}{2j\tau_{\rm m}},$ (19a)
 $\varphi_{\rm eq}$ $\simeq$ $\frac{2}{\tau_{e}}\Phi_{\rm eq}^{1/2},$ (19b)
 $\eta_{\rm eq}$ $\simeq$ $\frac{1}{2\sqrt{\Phi_{\rm eq}}}-2\Phi_{\rm eq}.$ (19c)
-We then use the linear perturbation method to calculate these quantities ($\Phi{=}\Phi_{\rm eq}+\delta\Phi$ etc.). The corresponding linearized equations read
-${\begin{bmatrix}\delta\dot{\Phi}\\ \delta\dot{\varphi}\\ \delta\dot{\eta}\end{bmatrix}=\begin{bmatrix}-\tau_{e}^{-1}&\Phi_{\rm eq}^{1/2}&0\\ -(0.25\Phi^{-3/2}+2)&-\tau_{e}^{-1}&-1\\ 4j\tau_{\rm e}^{-1}(j-1)^{-1}&0&0\end{bmatrix}\begin{bmatrix}\delta{\Phi}\\ \delta{\varphi}\\ \delta{\eta}\end{bmatrix}}.$ (20)
+We then use the linear perturbation method to analyze the system’s dynamics ($\Phi{=}\Phi_{\rm eq}+\delta\Phi$ etc.), and the corresponding linearized equations read
+[…]
 We assume that the solutions have the following forms $\delta\Phi,\delta\varphi,\delta\eta\propto e^{\lambda t}$ and $\lambda$ is a complex number. The characteristic equation can be written as
-$\lambda^{3}+\frac{2}{\tau_{e}}\lambda^{2}+\left(2\Phi_{\rm eq}^{1/2}+\frac{1}{4\Phi_{\rm eq}}\right)\lambda+\frac{\beta}{\tau_{a}\Phi_{\rm eq}^{1/2}}=0.$ (21)
+$\lambda^{3}+\frac{2}{\tau_{e}}\lambda^{2}+\left(2\Phi_{\rm eq}^{1/2}+\frac{1}{4\Phi_{\rm eq}}\right)\lambda+\frac{2\beta}{\tau_{\rm m}\Phi_{\rm eq}^{1/2}}=0.$ (21)
 We provide a general approach to solve the real part solutions of Eq.
-(21) as folows.
-Suppose that a cubic equation has a form of $A\lambda^{3}+B\lambda^{2}+C\lambda+D{=}0$ where $A,B,C,D{>}0$.
-Assuming that $\lambda{=}\lambda_{\rm r}\pm i\lambda_{\rm i}$, it can be proved that when $D{<}BC/A$, the real part of eigenvalues $\lambda_{r}{<}0$.
-This consequently means that the trapping is stable where the trajectory of Hamiltonian $H$ converges to the fixed point (see an illustration in Figure 2).
+(21) as follows.
+Suppose that a cubic equation has a form of $A\lambda^{3}+B\lambda^{2}+C\lambda+D{=}0$ where A,B,C,D>0A,B,C,D{>}0.
+Assuming that $\lambda{=}\lambda_{\rm r}\pm i\lambda_{\rm i}$, it can be proved that when D<BC/AD{<}BC/A, the real part of eigenvalues $\lambda_{r}{<}0$.
+This consequently means that the trapping is stable where the trajectory of Hamiltonian $H$ converges to the equilibrium (see an illustration in Figure 2).
 Substituting $A,B,C,D$ with Eq.
-(21), we obtain our stability criterion as
-$\frac{\tau_{a}}{\tau_{e}}>\left(\frac{3}{f_{\rm d,i}}\right)^{2/3}\frac{(j-1)^{2/9}j^{4/9}}{2j}\mu_{\rm o}^{-2/3}\sim\mu_{\rm o}^{-2/3}.$ (22)
+(21), we obtain the stability criterion as
+$\frac{\tau_{\rm m}}{\tau_{e}}>\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}\frac{1}{j}\left[\frac{(j-1)}{f_{\rm d,i}\alpha}\right]^{2/3}\sim\mu_{\rm o}^{-2/3}.$ (22)
 Our result is consistent with Eq. (26) of Goldreich & Schlichting (2014), where they constructed different forms of conserved quantities for derivations (see their Eq. 9).
 Our analysis can also be extended to the case when the mass of the outer planet is much lower than that of the inner one ($m_{\rm i}{\gg}m_{\rm o}$, $q{\gg}1$).
-In such a circumstance, we find that the real part of eigenvalues of characteristic equation $\lambda_{\rm r}{<}0$ always holds.
+In such a circumstance, we find that the real part of eigenvalues of the characteristic equation $\lambda_{\rm r}{<}0$ always holds.
 Importantly, this means that once the planet pair gets trapped in the resonance, the trapping is always stable as long as $m_{\rm i}{\gg}m_{\rm o}$.
-Figure 4: Level curves of the Hamiltonian for different $\eta$, plotted in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$.
+This holds for the planet pairs with an equal eccentricity damping timescale.
+Detailed analysis for the planet pairs with different $\tau_{\rm e}$ can be found in Deck & Batygin (2015).
+Figure 4: Level curves of the Hamiltonian (Eq.
+14) for different $\eta$, plotted in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$.
 The black (red) dots mark the stable (unstable) fixed points, the thick black line in panel (b) marks the separatrix.
@@
 The color of the lines represents the values of the Hamiltonian.
-When considering the planet-disk interaction, during the convergent migration of the planet, H increases, which corresponds to a color shift from purple and green to yellow.
-When overstability occurs, the perturbations grow, causing the curve to expand, and H decreases, corresponding to a color shift from yellow and green to purple.
 2.2.2 Escape criterion
-When planets are in resonances and under the stable criterion, the resonant quantities ($\phi$,$\Phi$,$\eta$) remain fixed with a small amplitude liberation.
+When planets are in resonances and under the stable criterion the resonant quantities ($\phi$$\Phi$$\eta$) remain fixed with a small amplitude libration.
 Here we attempt to answer what occurs if the above stability criterion is not satisfied.
-We define such a state as overstablity.
-In this case, the perturbations grow exponentially over time (no longer in the linear regime).
-Whether the overstability leads planets directly escape or still stay in resonance can be further analyzed by the phase space topology.
+This state is commonly described as overstability (Goldreich & Schlichting, 2014; Deck & Batygin, 2015; Xu & Lai, 2017).
+In this case, the perturbations grow exponentially over time due to non-adiabatic effects (eccentricity damping in this case).
+Whether the overstability leads planets to escape directly or still stay in resonance can be further analyzed by the phase space topology.
 Figure 4 shows the Hamiltonian curves for different $\eta$ values in the phase space of the conjugate variables $X{=}\sqrt{\Phi}\cos{\varphi}$ and $Y{=}\sqrt{\Phi}\sin{\varphi}$.
-We note that $\eta$ represents the deviation from the exact resonance and the distance to the origin corresponds to the excited eccentricity strength ($\sqrt{X^{2}+Y^{2}}{=}\sqrt{\Phi}{\sim}e$).
+We note that $\eta$ represents the deviation from the exact resonance and the distance to the origin corresponds to the excited eccentricity ($\sqrt{X^{2}+Y^{2}}{=}\sqrt{\Phi}{\sim}e$).
 Importantly, the value of $\eta$ sets the number of fixed points, and therefore determines the subsequent dynamical evolution.
-The fixed points can be obtained from $\partial H/\partial\varphi{=}\partial H/\partial\Phi{=}0$ (also see Eq.8-135 of Murray & Dermott 1999).
-It can be shown that, for $\eta{>}-3/2$, there is only one stable fixed point (black dot in Figure 4), whereas for $\eta{<}-3/2$, there are two stable fixed points and one unstable point (red dot in Figure 4).
-In the overstable circumstances, the perturbation near the equilibrium grows significantly.
-When the equilibrium value $\eta_{\rm eq}{>}-3/2$, the system’s Hamiltonian curve expands as eccentricities excited during resonant capture.
+The fixed points can be obtained from $\partial H/\partial\varphi{=}\partial H/\partial\Phi{=}0$.
+It can be shown that (Ferraz-Mello, 2007; Batygin, 2015; Petit et al., 2017), for $\eta{>}-3/2$, there is only one stable fixed point (black dot in Figure 4), whereas for $\eta{<}-3/2$, there are two stable fixed points and one unstable point (red dot in Figure 4).
+In overstable circumstances, the perturbations near the equilibrium grow over time.
+When the equilibrium value $\eta_{\rm eq}{>}-3/2$, the system’s Hamiltonian curve expands as eccentricities are excited during resonant capture.
 However, the trajectory still bounds to the only fixed point.
-As a result, the system’s eccentricity, period ratio, and resonant angle librate near the equilibrium.
-Following Goldreich & Schlichting (2014) and Xu & Lai (2017), we refer to this state as an overstable trap, which is also equivalent to circumstance described as the limit cycle in Afkanpour et al.
+As a result, the system’s eccentricity, period ratio, and resonant angle librate around their equilibrium values.
+Following Goldreich & Schlichting (2014) and Xu & Lai (2017), we refer to this state as an overstable trap – although the equilibrium state is overstable and perturbations grow over time ($\delta\Phi\propto e^{\lambda_{\rm r}t}$, $\lambda_{\rm r}{>}0$), the system remains trapped with the resonant angle librating at a finite amplitude.
+This state is also termed as the limit cycle in Deck & Batygin (2015); Nesvorný et al.
+(2022); Afkanpour et al.
 (2024).
-One example of overstable trap can be found in Figure 3.
-The separatrix occurs when $\eta_{\rm eq}{<}-3/2$ and the topology completely changes, as shown in Figure 4.
+One example of an overstable trap can be found in Figure 3.
+The separatrix emerges when $\eta_{\rm eq}{<}-3/2$ and the topology completely changes, as shown in Figure 4.
 In this case, the system’s Hamiltonian curve originally sits near the right stable fixed point.
-As the curve expands, it cross the separatrix and enter the circulation zone centered around the left stable fixed point (from the right to the left black dot in Figure 4).
-After then, the curve continuously shrinks towards the left black fixed point.
+As the curve expands, it crosses the separatrix and enters the circulation zone centered around the left stable fixed point (from the right to the left black dot in Figure 4).
+After that, the curve continuously shrinks towards the left black fixed point.
 As such, the eccentricity decreases, and the weaker resonant interaction cannot keep the planet pair in resonance anymore.
@@
 One example of escape is shown in Figure 5.
-Based on Eq. (19c), $\eta_{\rm eq}{<}-3/2$ requires $\Phi_{\rm eq}{=}(j-1)\beta\tau_{\rm e}/(4j\tau_{a}){>}1$, which is corresponding to
-$\frac{\tau_{a}}{\tau_{e}}<\left(\frac{3}{f_{\rm d,i}}\right)^{2/3}\frac{(j-1)^{8/9}j^{4/9}}{8j}\mu_{\rm o}^{-2/3}\sim\mu_{\rm o}^{-2/3}.$ (23)
+Based on Eq. (19c), $\eta_{\rm eq}{<}-3/2$ requires $\Phi_{\rm eq}{=}(j-1)\beta\tau_{e}/(2j\tau_{\rm m}){>}1$, which is corresponding to
+$\frac{\tau_{\rm m}}{\tau_{e}}<\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}\frac{1}{4j}\left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}\sim\mu_{\rm o}^{-2/3}.$ (23)
 The above escape criterion agrees with Eq. (28) of Goldreich & Schlichting (2014).
-It is worth noting that this escape condition requires the system is already in the overstable regime.
+It is worth noting that this escape condition requires the system already to be in the overstable regime.
 In the case where the mass of the outer planet is much lower than that of the inner one ($q\gg 1$), the resonance is always stable.
 As such, the escape does not occur even when $\eta_{\rm eq}{<}-3/2$.
-Taking together, resonance trapping is always stable when the mass of the inner planet is much higher than that of the outer one ($q{\gg}1$).
+To summarise, resonance trapping is always stable when the mass of the inner planet is much higher than that of the outer one ($q{\gg}1$).
 However, when the mass of the outer planet is much higher than that of the inner one ($q{\ll}1$), the outcome can be classified into three configurations: stable trap, overstable trap, or escape.
 We further extend the analysis to arbitrary mass ratios $q$ and derive the corresponding criteria as follows:
-$\begin{cases}{\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}].\end{cases}$ (24)
+$\begin{cases}{\displaystyle\frac{\tau_{\rm m}}{\tau_{e}}>\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)}{f_{\rm d,i}\alpha}\right]^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vskip 2.84544pt\\ {\displaystyle\frac{\tau_{\rm m}}{\tau_{e}}<\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}],\end{cases}$ (24)
 where
-$h(q)$ $=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (25)
+$h(q)$ $=$ $\frac{1}{(1+q\sqrt{\alpha})}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (25)
 $f(q)$ $=$ $1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (26)
-Mention that $h(q)=\tau_{a}/\tau_{e}(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2})$, $f(q)=8(j-1)[1-(e_{\rm o,eq}/e_{\rm i,eq})^{2}]$.
-Eq.
-(24) reduces to Eqs.
-(22) and (23) in the limit $q\ll 1$.
-We show the detailed derivation in the Appendix B.
-Figure 5: Resonance crossing of a two-planet system with convergent migration.
-The initial starting from the outer-to-inner planet period ratio is $2.2$ (slightly beyond the $2$:$1$ MMR), and migration parameters are $\tau_{a}=2.5\times 10^{5}$ and $\tau_{a}/\tau_{e}=10^{2}$.
-The period ratio between the inner and outer planets exhibits small oscillations upon entering resonance and ultimately decreases further away from the 2:1 ratio.
-The eccentricity of the inner planet undergoes significant oscillations around the resonant equilibrium value after resonance is established, eventually damping to a very small value.
-The resonant angle of the inner planet maintains large oscillations upon entering resonance and eventually diverges.
-In phase space, the inner planet initially oscillates around the resonant fixed point, but the oscillation range increases until it is ultimately captured by the origin.
-Figure 6: The theoretical distribution of the four kinds of evolution near 2:1 MMR depending on different $\tau_{a}$ and $\tau_{a}/\tau_{e}$.
-When $\tau_{a}$ and $\tau_{a}/\tau_{e}$ are in the blue region, the two planets pass through the resonance because the eccentricity damping is too strong.
-When $\tau_{a}$ and $\tau_{a}/\tau_{e}$ are in the blown region, they can’t be trapped into the resonance since they migrate inward too fast.
-The green region indicates that the planets can maintain a stable resonance.
-The yellow area represents that although the resonance trapping is overstable, the planets can still maintain a 2:1 period ratio.
-The purple region indicates that the planets will eventually leave the 2:1 MMR position and migrate inward after overstability.
-The left panel shows the distribution when the outer planet is more massive than the inner one, and the right panel displays the distribution when the inner planet is more massive than the outer one.
+We note that $h(q)$ and $f(q)$ can also be written as $h(q){=}\tau_{\rm m}/\tau_{e}(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2})$, and $f(q){=}1-(e_{\rm o,eq}/e_{\rm i,eq})^{2}$.
+In the limit $q{\ll}1$, Eq.
+(24) is equivalent to Eqs.
+(22) and (23) .
+We show the detailed derivation in Appendix B.
+Figure 5: Escape from resonance in a two-planet system with convergent migration.
+The planet and disk parameters are $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $\tau_{\rm m}{=}5\times 10^{5}$ yr and $\tau_{\rm m}/\tau_{e}{=}200$.
+The system is temporarily trapped in a 22:11 MMR during $t{\sim}2{-3}\times 10^{4}$ yr, and the inner planet reaches the equilibrium eccentricity with growing amplitude.
+The system escapes from resonance afterward, with $e_{\rm i}$ quickly damping to a very low value.
+Similarly, $\varphi_{\rm i}$ features an increasing libration amplitude during the trap and eventually oscillates from 0 to $2\pi$.
+In phase space, the trajectory initially librates around the right fixed point (black dot).
+As the amplitude grows, it ultimately escapes from the fixed point and settles into the origin.
+See https://github.com/llh-astro/resonance/raw/main/gif/escape.gif
+Figure 6: Analytical predictions of two-planet systems undergoing convergent migration near the 22:11 MMR in the limits of $q{\ll}1$ (left) and $q{\gg}1$ (right).
+Blue region: planets directly bypass the MMR due to strong eccentricity damping.
+Brown region: convergent migration is too fast to allow resonance capture.
+Green region: planets get trapped in a stable resonance.
+Yellow region: despite overstability occurs, planets still get trapped in the resonance, with libration of finite amplitude around the equilibrium point.
+Purple region: after a temporary trap, planets eventually escape resonance and continue the inward migration.
+The illustrative examples of the above regimes can be found in Figure 151{-}5.
 2.3 A unified picture
-We summarize different outcomes considering the dynamics evolution of planet pairs that undergo convergent migration.
-The illustrated runs from the numerical simulations in each regime are also presented.
-  1.
-No trap (Figure 1): The system cannot reach the resonant equilibrium state and the planet pair directly bypasses the resonance.
-The planets’ eccentricities increase and their period ratio temporarily holds during the resonant passage.
-However, after rapid crossing the resonance, the eccentricity drops and the resonant angle varies from 00 to $2\pi$.
+We summarize the resonance capture outcomes for planet pairs that undergo convergent migration. Representative simulation results from each regime are also presented.
+  1. No trap (Figure 2): The planets directly bypasses the resonance. After crossing the resonance, the planets’ eccentricities drop and the resonant angles vary from 0 to $2\pi$.
   2.
-Stable trap (Figure 2): The planets are locked in resonance with statically stable orbits.
-The planets’ period ratio, eccentricities, and resonant angle remain at the equilibrium values after capture.
-Meanwhile, the trajectory of the system converges towards a fix point in the phase space.
+Stable trap (Figure 2): The planets get locked in resonance, and the system’s trajectory converges towards an equilibrium value in the phase space.
+The planets’ period ratio, eccentricities, and resonant angles remain at equilibrium values after capture.
   3.
-Overstable trap (Figure 3): The planets are locked in resonance with dynamically stable orbits.
-Their eccentricities and resonance angle initially converge to the equilibrium values.
-After a while, the eccentricities and resonance angle liberate with increasing amplitudes.
-This state is nevertheless stable and the period ratio of planets remains constant after capture.
-In the phase space, the system’s trajectory initially approaches the fixed point, then begins to oscillate, and ultimately settles into a limit cycle around that point.
+Overstable trap (Figure 3): The planets are locked in resonance with librating resonant angles.
+The eccentricities and the resonance angles of planets initially converge to the equilibrium values and, after a while, librate with increasing amplitudes.
+The period ratio of the planets remains constant after capture.
+In the phase space, the system’s trajectory initially approaches the fixed point, then transitions into oscillations before ultimately stabilizing into a limit cycle around that point.
   4.
-Escape (Figure 5): The planets are locked in resonance with only temporarily stable orbits.
-Their period ratio, eccentricities and resonance angle initially converge to the equilibrium values.
-The planets’ eccentricities and resonance angle quickly oscillate with increasing amplitudes.
-This state becomes unstable and the planets eventually leave resonance.
-As such, the planets’ eccentricities decrease, and the period ratio continues to drop, and the resonant angle changes from liberation to circulation.
-In the phase space, the system’s trajectory initially approaches the fixed point; however, the oscillation gradually grows until it intersects the separatrix.
-The trajectory ultimately settles towards the coordinate origin.
-We point out that although several literature studies have explored this issue in depth (Huang & Ormel, 2023; Goldreich & Schlichting, 2014; Deck & Batygin, 2015; Batygin, 2015), none have formulated a comprehensive model.
-Here we demonstrate a unified picture in Figure 6.
-The outcome is given in the $\tau_{a}$-$\tau_{e}$ parameter space.
-We describe how $\tau_{\rm a}$ and $\tau_{\rm e}$ influence the outcomes as follows.
-If convergent migration is too rapid, the planets directly bypass the resonance without reaching the equilibrium state.
+Escape (Figure 5): The planets are temporarily locked in resonance before eventually escaping.
+The period ratio, eccentricities, and resonance angle of the planet pair initially converge to the equilibrium values.
+Then the eccentricities and resonance angles start to oscillate with growing amplitudes.
+Eventually, the planets leave resonance, with a shift in resonant angles from libration to circulation.
+In phase space, the system’s trajectory initially approaches the fixed point; however, the oscillation gradually amplifies until it crosses the separatrix.
+The trajectory ultimately stabilizes towards the coordinate origin.
+We point out that several previous studies have explored this issue in depth (Huang & Ormel, 2023; Goldreich & Schlichting, 2014; Deck & Batygin, 2015; Batygin, 2015; Pichierri et al., 2018; Batygin & Petit, 2023).
+Nevertheless, most studies focus on one aspect of these issues at a time.
+Here, we provide a unified framework for studying the dynamics of both resonance capture and subsequent stability.
+This analysis is applicable to planet pair systems with arbitrary planet mass ratios and at different first-order mean motion resonances.
+An illustration of this framework can be found in Figure 6.
+The results are presented in a $\tau_{\rm m}/\tau_{e}$$\tau_{\rm e}$ parameter space, which links all dynamical regimes to the two key parameters that govern planet-disk interactions: angular momentum and eccentricity damping.
+We describe below how the variations in $\tau_{\rm m}$ and $\tau_{\rm e}$ influence the outcome.
+If convergent migration occurs too rapidly, the planets directly bypass the resonance without reaching the equilibrium state.
 As migration slows down, the effect of eccentricity damping becomes significant.
-Strong e-damping has two consequences: the planet resonant interaction gets weaken and semimajor axis decays due to orbital circularization.
-When the orbital decay surpasses the resonant term, the planets will no longer remain in resonance.
-In this regard, planets cannot get capture when $\tau_{\rm e}$ and/or $\tau_{\rm a}$ are short.
-No trap occurs in the upper and left part of Figure 6.
-When planets are trapped in resonance, the stability is related to the degree of eccentricity excitation.
-The damping effect ($\tau_{\rm e}$) reduces the planet’s eccentricity, leading the system stable.
-On the other hand, convergent migration ($\tau_{\rm a}$) brings planets closer together, in turn exciting their eccentricities.
-Consequently, the relative eccentricity damping $\tau_{\rm a}/\tau_{\rm e}$ is the key parameter affecting the stability.
-As can be seen from top to bottom of Figure 6, system tends to be more unstable as $\tau_{\rm a}/\tau_{\rm e}$ decreases.
-For the small inner planet case shown in Figure 6, if the relative damping $\tau_{\rm a}/\tau_{\rm e}$ is sufficiently strong, the resonant trapping is statically stable, and the final state of the system stays at a fixed point in the phase space (stable trap regime).
+Strong e-damping has two effects.
+First, it causes a reduction in planet’s eccentricity, weakening the resonant interaction, since the resonant term is tied to planet’s eccentricity ((3)).
+Furthermore, the semi-major axis also decays due to orbital circularization.
+If the magnitude of such orbital decay term surpasses the resonant term, the planets will no longer remain in resonance.
+In this regard, planets cannot be captured when $\tau_{e}$ and/or $\tau_{\rm m}$ are short.
+No trap occurs in the upper and left parts of Figure 6.
+When planets are trapped in resonance, the long-term stability of the system is related to the degree of eccentricity excitation.
+The damping effect ($\tau_{e}$) reduces the planet’s eccentricity, making the system stable.
+On the other hand, convergent migration ($\tau_{\rm m}$) brings planets closer together, in turn exciting their eccentricities.
+Consequently, the relative eccentricity damping $\tau_{\rm m}/\tau_{e}$ is the key parameter in determining the stability.
+As can be seen from the top to the bottom of Figure 6, the system tends to be more unstable as $\tau_{\rm m}/\tau_{e}$ decreases.
+For the small inner planet case shown in Figure 6, if the relative damping $\tau_{\rm m}/\tau_{e}$ is sufficiently strong, the resonant trapping is statically stable, and the final state of the system stays at a fixed point in the phase space (stable trap regime).
 As the relative damping weakens (Eq.
-22), the fixed point becomes unstable, and the system’s trajectory begins to oscillate around the fixed point.
+22), the fixed point becomes unstable and the trajectory of the system begins to oscillate around the fixed point.
 It ultimately settles into a stable limit cycle (overstable trap).
-The limit cycle expands as $\tau_{\rm a}/\tau_{\rm e}$ continues to decrease.
+The limit cycle expands as $\tau_{\rm m}/\tau_{e}$ continues to decrease.
 When the trajectory crosses the separatrix (Eq.
-23), the system becomes dynamically unstable.
+23), the system becomes unstable.
 In this regime, the oscillations grow over time and planets eventually leave resonance (escape).
 In contrast, for the case of a small outer planet shown in Figure 6, all traps are stable.
-Figure 7: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q$ of $0.1$.
-Blue points indicate cases where the planets are not trapped by the resonance, green points indicate stable resonance capture, yellow points indicate overstable resonance capture, and purple points indicate that the planets eventually escape from the resonance.
-The simulations generally align with the schematic shown in Figure 6, except for the region near $\tau_{a}\sim 10^{5}\ \mathrm{yr}$ and $\tau_{a}/\tau_{e}\sim 3\times 10^{2}$.
-While the analytical model predicts overstable trapping in this regime, the simulations show escape from resonance.
-We hypothesize that as migration becomes faster, the perturbations grow stronger and become nonlinear, leading to amplified oscillation amplitudes and an increased likelihood of escaping from resonance.
-Figure 8: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q$ of $10$.
-The simulations generally align with the schematic shown in Figure 6, except for the region near $\tau_{\rm a}{\sim}2{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm a}/\tau_{\rm e}{\sim}2\times 10^{2}$.
-While the analytical model predicts resonance trapping in this regime, the simulations show resonance crossing.
-We hypothesize that the discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect.
-Figure 9: Numerical results of 2:1 MMR capture and stability in $\tau_{a}-\tau_{a}/\tau_{e}$ parameter space with inner to outer planetary mass ratio $q=1$.
-We compare our criteria with previous studies.We cite the trapping criterion from Huang & Ormel (2023) and the stable and escape criteria from Goldreich & Schlichting (2014).
-The solid lines represent our criteria, the dash lines represent the criteria cited from previous studies.
-In the case of equal-mass planets, it is evident that our criteria align more closely with the simulation results.
-3 Simulations
-In this section, we conduct numerical simulations for $2:1$ resonance to verify the different regimes and criteria analytically derived in the previous section.
-The numerical setup is described in Sec.
-3.1.
-The comparison between simulations and analysis at two extreme mass ratios is presented in Sec.
-3.2, and the extrapolation for any mass ratio is discussed in Sec.
-3.3.
-3.1 Numerical setup
-We use the open-source N-body code REBOUND with the WHfast integrator (Rein & Liu, 2012).
-The simulations are confined in conplanar circumstances, and the accelerations for orbital decay and eccentricity damping are given by
-${a_{\rm m}}=-\frac{\textbf{v}}{2\tau_{a}},\ \ {a_{\rm e}}=-2\frac{(\textbf{v}\cdot\textbf{r})\cdot\textbf{r}}{r^{2}\tau_{e}},$ (27)
-which are implemented through REBOUNDx (Tamayo et al., 2020).
-Initially, we place the inner and outer planets at $1$ and $1.7$ au, respectively, making them slightly away from the $2$:$1$ MMR.
-Both planets have circular orbits with random phase angles.
-To mimic convergent migration, we only let the outer planet undergo inward migration on a time-scale of $\tau_{\rm a}$, whereas the eccentricities of both planets are damped on a time-scale of $\tau_{\rm e}$.
-Simulations are terminated at $t{=}2\tau_{\rm a}$, and the final outcomes of the planets in each run are determined as described in Sec.
-2.3.
-We consider three cases for the masses of the planet-pair: small inner planet ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $q{=}0.1$), small outer planet ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, $q{=}10$), equal-mass planets ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, $q{=}1$).
-We vary different sets of $\tau_{\rm a}$ and $\tau_{\rm e}$ in each simulations.
-3.2 Comparison at two extreme mass ratios
-The dynamical outcomes of planet pairs that undergo convergent migration are shown in Figure 6.
-The results are dependent on the mass ratio of the planets and disk properties ($\tau_{\rm a}$ and $\tau_{\rm e}$).
-Basically four regimes are presented when the inner planet is lower than that of the outer one.
-The illustrated runs in each regime are provided in Figures 1,2,3 and 5.
-Figure 8 shows the numerical results obtained from an illustrated case where the mass of inner planet is ten times lower than that of the outer one ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, $q{=}0.1$).
-The numerical results shown in Figure 8 can directly be compared with the schematic presented in Figure 6.
-We find that the simulations agree pretty well with the analytical estimates.
-No planets lead resonant trap (blue dots) in the upper left part of Figure 8, when the weak damping (upon the green line) and slow migration (left of the brown line) criteria break.
-When those trapping criteria are satisfied and as $\tau_{\rm a}/\tau_{\rm e}$ decreases, the simulations depict a transition from stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes.
-These results are indeed separated by the two analytical criteria proposed in Sect.
-2 (yellow and purple lines).
-However, there is one point in Figure 8 that simulations do not fully match the analysis, occurring at $\tau_{a}{\sim}10^{5}\ {\rm yr}$ and $\tau_{a}/\tau_{e}{\sim}3\times 10^{2}$.
-The theory predicts the planets in overstable trap regime whereas in simulations planets actually leave the resonance.
-We speculate that the reason is that as migration becomes faster, the perturbations intensify and become nonlinear, resulting in further increasing oscillation amplitudes around the equilibrium value.
-Consequently, the system becomes more likely to cross the separatrix and escape resonance.
-Figure 8 demonstrates simulation results from the case when the mass of inner planet is ten times higher than that of the outer one ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, $q{=}10$), serving as a comparison to the schematic illustrated in Figure 6.
-Again, no trap (blue dots) occurs in the upper and left parts of Figure 8, when the weak damping and slow migration criteria are not satisfied.
-When these criteria are fulfilled, all traps correspond to stable trap (green dots).
-We note that the no-trap regime occupies a large parameter space than theoretically expected.
-This is evident at $\tau_{\rm a}{\sim}2{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm a}/\tau_{\rm e}{\sim}2\times 10^{2}$, where the trapping criteria are fulfilled but the planets are not in resonance.
-This discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect (Eq.
-6).
-When the eccentricity damping is considered, the interaction between the planets is further suppressed (lower $e$ in left-hand side of Eq.
-3), leading to a longer migration timescale (right-hand side of Eq.
-3).
-Consequently, in reality planets can bypass the resonance with slower migration rates compared to the simplified estimation from Eq.
-6.
-In conclusion, our analytical predictions align well with simulations in the circumstances of planet pairs with extreme mass ratios.
-3.3 Comparison at arbitrary mass ratios
-Notably, Huang & Ormel (2023) derived the weak damping trapping criterion at the small mass ratio limit of $q{\ll}1$.
-Goldreich & Schlichting (2014) investigated when the trap becomes overstability and/or escape in the same limit.
-In this work we derived the general criteria for trap and stability that are well suited to the arbitrary mass ratio (Eqs.
-13 and 24).
-In this subsection we show the simulation comparison, whereas the detailed derivation of the $q$-dependent overstabity and escape criteria is provided in Appendix B.
-Figure 9 illustrates the simulation results for the equal-mass planet case ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, $q{=}1$) and compares our criteria with those from Huang & Ormel (2023) and Goldreich & Schlichting (2014).
-The solid lines represent our analytical criteria, dashed lines correspond to criteria from previous studies, and colored dots depict the simulation results.
-We find that the new criteria clearly depict a better agreement with numerical results.
-Figure 10: The plots illustrate the dependence of resonance trapping criteria on the semi-major axis (panel a) and the stellar mass (panel b) for a two-planet system, based on the minimum solar nebular model (MMSN).
-The planet masses are $m_{\rm i}=1\ M_{\oplus}$ and $m_{\rm o}=10\ M_{\oplus}$.
-In panel a, the stellar mass is fixed at $M_{\star}=1\ M_{\odot}$, while in panel b, the semi-major axis of the inner planet is fixed at $a_{\rm i}=1\ au$.
-The brown line represents the slow migration trapping criterion (Eq.
-8), below which migration is too fast for resonance capture.
-Regions above the brown line indicate where capture is possible.
-The color gradient represents the strongest eccentricity damping $\tau_{a0}/\tau_{e0}$ that allows resonant capture (Eq.
-11), with lighter colors meaning the planets can be in resonances with stronger eccentricity damping.
-In panel (a), as the semi-major axis increases, the critical value of $\tau_{a0}$ remains constant, while $\tau_{a0}/\tau_{e0}$ slightly increases.
-This suggests that planets are more easily trapped in resonances under stronger eccentricity damping as the semi-major axis increases.
-In panel (b), as the stellar mass decreases, the critical value of $\tau_{a0}$ decreases, and $\tau_{a0}/\tau_{e0}$ increases.
-This indicates that planets are more easily trapped in resonances under stronger eccentricity damping and faster migration as the stellar mass decreases.
-4 Discussion
-In this section, we connect our criteria to disk parameters in Sec.
-4.1 and analyze how the criteria depend on the planet location and the mass of stellar host.
-Additionally, we investigate resonance trapping in the context of divergent migration in Sec.
-4.2 and discuss several caveats of our work in Sec.
-4.3.
-4.1 Dependence on semi-major axis and stellar mass
-Here we explore how the resonant trapping conditions depend on the radial distance and stellar mass.
-The detailed type I migration formulas for low-mass planets are given by $\tau_{a}^{-1}{\sim}\mu\mu_{\rm d}h^{-2}n$ and $\tau_{e}^{-1}\sim\mu\mu_{\rm d}h^{-4}n$, where $\mu_{\rm d}{=}\Sigma_{\rm d}a^{2}/M_{\rm*}$ is the mass ratio between disk and star, and $\Sigma_{\rm d}$ and $h$ are the disk surface density and aspect ratio, respectively.
-Following the minimum mass solar nebula (MMSN, Hayashi 1981) model, the disk surface density and aspect ratio can be written as
-$\Sigma_{\rm d}=$ $1700\left(\frac{a}{\rm 1\ au}\right)^{-3/2}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)\rm\ g/cm^{2},$ (28)
-$h=$ $0.033\left(\frac{a}{\rm 1\ au}\right)^{1/4}\left(\frac{L_{\star}}{1\ L_{\odot}}\right)^{1/8}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{-1/2}$
-$=$ $0.033\left(\frac{a}{\rm 1\ au}\right)^{1/4}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{-1/4},$ (29)
-the latter of which $L_{\star}{\propto}M_{\star}^{2}$ is assumed for the pre-main sequence stars.
-By substituting $\Sigma_{\rm d}$ and $h$ into $\tau_{a}$ and $\tau_{e}$, we obtain their dependence on the semi-major axis and stellar mass as
-$\tau_{a}={\tau_{a0}}\left(\frac{a}{\rm 1\ au}\right)^{3/2},\ \ \tau_{e}={\tau_{e0}}\left(\frac{a}{\rm 1\ au}\right)^{2}\left(\frac{M_{\star}}{1\ M_{\odot}}\right)^{1/2},$ (30)
-where $\tau_{a0}$ and $\tau_{e0}$ represent the values of $\tau_{a}$ and $\tau_{e}$ at $1$ au around a solar-mass host.
-By inserting $\tau_{a0}$ and $\tau_{e0}$ into Eq.
-(13), we can separate the dependence of resonance trapping on semi-major axis and stellar mass.
-Figure 10 shows the a-dependency for resonance trapping, where $M_{\star}{=}1\ M_{\odot}$ and $m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$.
-The region below the brown line (slow migration criterion) indicates that migration is too fast for resonance capture, while the region above suggests that capture is possible, which depends on $\tau_{a0}/\tau_{e0}$.
-The color in Figure 10 refers to the critical $\left(\tau_{a0}/\tau_{e0}\right)$ that allows resonant capture.
-A lighter color means the planets can be in resonances with higher $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ (stronger e-damping).
-As the semi-major axis increases, $(\tau_{a0})_{\rm slow,mig}$ remains constant and $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ slightly increases.
-This means that planets would be more easily trapped at resonances under stronger e-damping as $a$ increases.
-Nevertheless, the capture reveals an overall weak semi-major axis dependence.
-Figure 10 shows $M_{\star}$-dependency for resonance trapping.
-We fix the inner planet at $1\ au$ and vary with the stellar mass.
-From the figure, as the stellar mass decreases, $(\tau_{a0})_{\rm slow,mig}$ decreases, and $\left(\tau_{a0}/\tau_{e0}\right)_{\rm weak,e}$ increases.
-This implies that planets can be trapped under a wider range of $\tau_{a}$ and $\tau_{e}$ parameter space.
-In other words, planets are more likely to be capture into resonance as the stellar mass decreases.
-Based on the MMSN model, we find that planets are more likely to get capture into resonance with increasing semi-major axis and/or decreasing stellar mass.
-However, we note that these results are considerably sensitive to the adopted disk model.
-The interpretation should be taken with caution.
-4.2 Divergent migration
-We have shown in the previous sections that the convergent planet migration coupled with eccentricity damping can result in resonance trapping.
-Literature theoretical work suggested that planets cannot be captured in resonance when the direction of their migration is divergent (Wong & Lee, 2024).
-Based on our analysis, here we also provide a simple argument that why resonance trapping cannot occur through divergent migration.
+Figure 7: Simulations of the 22:11 MMR capture and stability for two-planet systems with an inner to outer planet mass ratio $q$ of $0.1$.
+Blue: planets bypass resonance due to rapid migration or strong eccentricity damping.
+Green: stable resonance capture.
+Yellow: overstable resonant capture.
+Purple: Resonance escape after temporary capture.
+The simulations are consistent with the analysis outlined in Figure 6.
+Figure 8: Same as Figure 7 but with an inner to outer planet mass ratio $q$ of $10$.
+Figure 9: Same as Figure 7 but with equal-mass of q=1q{=}1.
+The solid lines are the criteria derived in this study, whereas the dashed lines represent the trapping criteria of Batygin (2015) (their Eq.
+43), and the stability and escape criteria of Deck & Batygin (2015) (their Eqs.
+15 and 21).
+2.4 Resonance crossing through divergent migration
+We have shown in the previous subsections that the convergent planet migration coupled with eccentricity damping can result in resonance trapping.
+Literature theoretical work suggested that planets directly cross the resonance when the direction of their migration is divergent due to separatrix crossing (Henrard & Lemaitre, 1983; Murray & Dermott, 1999; Wong & Lee, 2024).
+Here we also explain that divergent migration cannot lead to resonance capture in our framework.
 When the planet pair undergoes the divergent migration, the equilibrium state $\dot{n_{\rm i}}/n_{\rm i}{=}\dot{n_{\rm o}}/n_{\rm o}$ (if it exists) can be written as
 $-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})$ (27)
-$=-\frac{1}{2\tau_{a}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
-where the relative orbital decay term ($-\tau_{a}^{-1}$) is negative compared to Eq.
+$=-\frac{1}{\tau_{\rm m}}+\frac{e_{\rm o}^{2}-e_{\rm i}^{2}}{\tau_{e}},$
+where the relative orbital decay term ($\tau_{\rm m}^{-1}$) is negative compared to Eq.
 (3).
@@
 (27) reads
-$\frac{e_{\rm i}^{2}}{\tau_{\rm e}}(1+q\sqrt{\alpha})[j+(j-1)\frac{f_{\rm d,o}^{{}^{\prime}2}}{f_{\rm d,i}^{2}}q\sqrt{\alpha}]=-\frac{1}{2\tau_{a}}.$ (32)
-Noticeably, the left-hand side of this equation is positive, whereas the right-hand side is negative.
-This indicates that there is no equilibrium solution in the divergent migration circumstance.
-We also conducted numerical simulations and consistently found that even a small amplitude of divergent migration is sufficient for planets to cross the resonance.
-4.3 Caveats
-Here discuss a few limitations in our model.
-First, for simplicity we treat $\tau_{\rm a}$ and $\tau_{\rm e}$ as free but constant parameters and two planets have the same eccentricity damping timescales.
-However, in reality $\tau_{\rm a}$ and $\tau_{\rm e}$ depend on the gas disk mass, planet mass and orbital distance.
-For instance, $\tau_{\rm a}$ and $\tau_{\rm e}$ decrease as planet mass increases, which scale with $\mu^{-1}$.
-The two trapping criteria exhibit steeper decreasing functions of $\mu$ such that $\tau_{\rm a,\ \text{slow\ mig}}{\propto}\mu^{-4/3}$, and $(\tau_{\rm a}\tau_{\rm e})_{\rm\text{weak\ damp}}{\propto}\mu^{-2}$.
-This suggests that planets become more easily trapped as their mass increases.
-On the other hand, since the planets can grow their masses at different rates, the planet mass ratio could be time varied.
-For instance, if the inner planet is initially more massive than the outer planet ($q{\gg}1$), the resonance trapping is the stable regime.
+$\frac{e_{\rm i}^{2}}{\tau_{\rm e}}(1+q\sqrt{\alpha})[j+(j-1)\frac{f_{\rm d,o}^{{}^{\prime}2}}{f_{\rm d,i}^{2}}q\sqrt{\alpha}]=-\frac{1}{\tau_{\rm m}}.$ (28)
+Noticeably, as long as $\tau_{\rm m}$ and $\tau_{\rm e}$ have the same sign, the left-hand side of this equation has the opposite sign as the right-hand side, and therefore the equality never holds.
+This indicates that there is no equilibrium solution in the divergent migration circumstance, and even a small amplitude of divergent migration is sufficient for planets to cross the resonance.
+3 Simulations
+In this section, we conduct numerical simulations to verify the aforementioned analytical criteria.
+The numerical setup is described in Section 3.1.
+We first test the 22:11 resonance for planets with extreme mass ratios in Section 3.2, and then extend the analysis for planets with arbitrary mass ratios in Section 3.3.
+Other first-order resonances are investigated in Section 3.4.
+3.1 Numerical setup
+We use the open-source N-body code REBOUND with the WHfast integrator (Rein & Liu, 2012).
+In simulations we also consider planets in coplanar orbits.
+The accelerations for angular momentum and eccentricity damping are given by (Papaloizou & Larwood, 2000; Cresswell & Nelson, 2006, 2008; Pichierri et al., 2024)
+${a_{\rm m}}=-\frac{\textbf{v}_{\rm pl}}{\tau_{\rm m}},\ \ {a_{\rm e}}=-2\frac{(\textbf{v}_{\rm pl}\cdot\textbf{r}_{\rm pl})\cdot\textbf{r}_{\rm pl}}{r_{\rm pl}^{2}\tau_{e}},$ (29)
+which are implemented through REBOUNDx (Tamayo et al., 2020).
+Initially, the inner planet is placed at $1$ au and the orbit of the outer planet is slightly away from the desired first-order resonance ( e.g., at $1.7$ au for the 22:11 resonance).
+Both planets have circular orbits with random phase angles.
+To mimic convergent migration, only the outer planet undergoes inward migration on a timescale of $\tau_{\rm m}$, and the eccentricities of both planets are damped on a timescale of $\tau_{e}$.
+Simulations are terminated at $t{=}2\tau_{\rm m}$, and the outcomes can be classified in different regimes, as described in Section 2.3.
+We consider three cases for the masses of the planet-pair: small inner planet ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, q=0.1q{=}0.1), small outer planet ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, q=10q{=}10), and equal-mass planets ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, q=1q{=}1).
+We vary $\tau_{\rm m}$ and $\tau_{e}$ in each simulation.
+3.2 Systems at two extreme mass ratios
+Figure 8 shows the simulations for the planet pair with the inner planet mass ten times lower than that of the outer one ($m_{\rm i}{=}1\ M_{\oplus}$, $m_{\rm o}{=}10\ M_{\oplus}$, q=0.1q{=}0.1), the results of which can be compared with the schematic presented in Figure 6.
+We find that the simulations agree pretty well with the analytical estimates.
+No planets lead the resonant trap (blue dots) in the upper left part of Figure 8, when the weak damping (green line) and slow migration (brown line) criteria break.
+When those trapping criteria are satisfied and as $\tau_{\rm m}/\tau_{e}$ decreases, the simulations depict a transition from the stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes.
+These results are indeed separated by the two analytical criteria proposed in Sect.
+2 (yellow and purple lines).
+The simulations do not fully match the analysis at $\tau_{\rm m}{\sim}2\ \times 10^{5}\ {\rm yr}$ and $\tau_{\rm m}/\tau_{e}{\sim}6\times 10^{2}$ in Figure 8.
+The theory predicts the planets in an overstable trap regime whereas in simulations planets actually leave the resonance.
+We hypothesize that the increased relative migration rate leads to more intense and nonlinear perturbations.
+This results in greater oscillation amplitudes around the equilibrium point.
+As such, the system has a higher probability of crossing the separatrix and escaping resonance.
+Figure 8 demonstrates simulations for the planet pair with the inner planet mass ten times higher than that of the outer one ($m_{\rm i}{=}10\ M_{\oplus}$, $m_{\rm o}{=}1\ M_{\oplus}$, q=10q{=}10), offering a comparison to the schematic illustrated in Figure 6.
+Again, no trap (blue dots) occurs in the upper and left parts of Figure 8, when the weak damping and slow migration criteria are not satisfied.
+When these criteria are fulfilled, all traps correspond to stable traps (green dots).
+We note that the no-trap regime occupies a slightly larger parameter space than theoretically expected.
+For instance, at $\tau_{\rm m}{\sim}4{\times}10^{5}\ {\rm yr}$ and $\tau_{\rm m}/\tau_{e}{\sim}4\times 10^{2}$, while the trapping criteria are fulfilled, the planets are not in resonance.
+This discrepancy arises from the simplification in deriving our slow migration criterion, which neglects the eccentricity damping effect (Eq.
+6).
+When the eccentricity damping is incorporated, the interaction between the planets is further suppressed—owing to reduced eccentricities in the left-hand side of Eq.
+3—thereby prolonging the migration timescale on the right-hand side of Eq.
+3.
+Consequently, planets in reality bypass the resonance with slower migration rates compared to the simplified estimation from Eq.
+6.
+Overall, we find that the analytical criteria align well with simulations in the circumstances of the planet pairs with extreme mass ratios.
+3.3 systems at arbitrary mass ratios
+The resonance dynamics has been extensively investigated in literature studies.
+Batygin (2015) proposed an adiabatic capture criterion for systems without eccentricity damping, linking the resonant libration time to the migration time.
+Subsequently, Huang & Ormel (2023) derived a weak eccentricity damping criterion for planet pairs with extreme mass ratios ($q{\ll}1$), while Batygin & Petit (2023) formulated a generalized criterion applicable to arbitrary mass ratios.
+Regarding the post-capture evolution, Goldreich & Schlichting (2014) first proposed the concept of overstability and derived the corresponding analytic criteria for restricted three-body systems.
+Deck & Batygin (2015) further extended the analysis for planet pairs with arbitrary mass ratios.
+In this subsection, we conduct numerical simulations to validate our theoretical criteria and compare our results with those from literature studies.
+The detailed derivation of the $q$-dependent overstability and escape criteria are provided in Appendix B.
+As a point of comparison, our weak damping criterion is equivalent to that obtained by Batygin & Petit (2023), while our other criteria exhibit the same $\mu$-scaling as those in previous works but differ in their numerical prefactors: $(\tau_{\rm m})_{\rm slow,mig}{\sim}\mu^{-4/3}$, $(\tau_{\rm m}/\tau_{e})_{\rm overstab},(\tau_{\rm m}/\tau_{e})_{\rm esc}{\sim}\mu^{-2/3}$.
+Therefore, we compare our slow migration criterion with the adiabatic criterion in Batygin (2015) and our stability and escape criteria with those in Deck & Batygin (2015).
+Additionally, we perform simulations to investigate the $q$-dependence of our overstability and escape criteria, as detailed in Appendix B.
+Figure 9 illustrates the simulations for the planet pair of equal-mass ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$, q=1q{=}1).
+The colored symbol depicts the simulation results, the solid lines represent our analytical criteria (Eq.
+24), and the dashed lines correspond to the criteria from Batygin (2015) (their Eq 43) and Deck & Batygin (2015) (their Eqs 15 and 21).
+As can be seen, both the criteria derived in this work and those from previous studies match the numerical results relatively well.
+A more detailed comparison that accounts for the dependence on the mass ratio $q$ is presented in Appendix B.
+3.4 Systems at arbitrary j:j-1 MMR
+In previous sections we focus on the circumstances of the 22:11 resonance as a representative case.
+Since our analytical framework can be applied to other first-order $j$:j1j-1 MMRs, we also perform such numerical verification as follows.
+Figure 10 shows the resonant states in which the systems finally settle for equal-mass planet pairs ($m_{\rm i}{=}m_{\rm o}{=}10\ M_{\oplus}$).
+The colored dots depict the numerical results, the green lines represent the trapping criteria, while the purple line represents the escape criterion.
+Note that for equal-mass planet pairs, our criteria predict that all traps are stable (no overstable trap and escape regimes) at $j{\geq}3$ (Eq.
+24).
+Therefore, we only exhibit the escape criterion for the 22:11 MMR in Figure 10.
+Our criteria show good agreement with simulations.
+As the disk effects become stronger (e.g., $\tau_{\rm m}$ decreases and $\tau_{\rm m}/\tau_{e}$ increases), planets can bypass the lower $j$:j1j-1 MMRs and evolve into more compact orbits.
+Figure 10: Simulations of $j$:j1j-1 MMR capture and stability for two-planet systems with equal masses.
+Color dots represent systems’ final resonant states.
+Green lines show $j$:j1j-1 trapping criteria, while purple indicates the 22:11 escape criterion.
+4 Discussion
+In this section, we outline a few limitations in our model.
+First, for simplicity we treat $\tau_{\rm m}$ and $\tau_{e}$ as free but constant parameters and two planets have the same eccentricity damping timescales.
+However, in reality $\tau_{\rm m}$ and $\tau_{e}$ depend on the gas disk mass, the planet mass, and orbital distance.
+For instance, in the linear type I regime, both $\tau_{\rm m}$ and $\tau_{e}{\propto}\mu^{-1}$, while the two trapping criteria exhibit steeper decreasing functions of $\mu$ such that $\tau_{\rm m,\ \text{slow\ mig}}{\propto}\mu^{-4/3}$, and $(\tau_{\rm m}\tau_{e})_{\rm\text{weak\ damp}}{\propto}\mu^{-2}$.
+This suggests that planets become more easily trapped as their masses increase.
+Nevertheless, we note that the relation between planet mass and trapping outcome behavors more complicated.
+When the planets’ masses further increase and exceed the gap-opening threshold, they gradually switch to Type-II migration (Lin & Papaloizou, 1986; Kanagawa et al., 2018).
+In this regime, the planet’s mass influences migration and eccentricity damping in a non-trivial way (Kanagawa et al., 2018; Kajtazi et al., 2023; Pichierri et al., 2023, 2024).
+This additional complexity is not accounted for in our current analysis.
+On the other hand, since the planets can grow their masses at different rates, the planet mass ratio could change over time.
+For instance, if the inner planet is initially more massive than the outer planet ($q{\gg}1$), the resonance trapping is in the stable regime.
 When the outer planet grows and the mass ratio decreases, the system may transition into an overstable trap or even escape.
-Moreover, both hydrodynamic simulations and analytical work have shown that the migration and eccentricity damping rates differ from those in the near circular case when planets attain high eccentricities (Cresswell & Nelson, 2008; Ida et al., 2020).
-Future studies are required to investigate how these factors influence our analysis.
-Second, we only consider a static disk.
-The configuration may change when disk dissipation is taken into account.
-In this case, the migration slows down as the gas density drops, making the resonance trapping easier.
+Future studies should explore how these interactions modify the trapping criteria and how the evolving planet mass ratio affects the long-term dynamics of the system.
+Furthermore, both hydrodynamic simulations and analytical work have shown that the migration and eccentricity damping rates differ from those in the near circular case when planets attain high eccentricities (Cresswell & Nelson, 2008; Bitsch & Kley, 2011; Bitsch et al., 2013; Xu et al., 2018; Ida et al., 2020; Pichierri et al., 2023, 2024; Fairbairn & Rafikov, 2025).
+For instance, Xu et al.
+(2018) showed that nonlinear eccentricity damping can lead to higher equilibrium eccentricities, enhance the stability of captured resonances.
+In addition, our analysis carried out so far assumes a laminar disk.
+In a more realistic turbulent disk, turbulence-induced density fluctuations can generate additional stochastic torques (Laughlin et al., 2004; Ogihara et al., 2006; Rein, 2012; Okuzumi & Ormel, 2013; Hühn et al., 2021; Wu et al., 2024).
+These effects are not accounted for in the aforementioned simplified framework.
+Such stochastic torques may considerably alter the resonance trapping outcomes (Batygin & Adams, 2017; Goldberg & Batygin, 2023; Chen et al., 2025).
+Lastly, we only consider a static disk without considering the gas density decay due to disk dissipation.
+As disk gas disperses, the planet migration slows down and resonance trapping is easier to occur.
 However, for stability/overstability and escape criteria, the situation is less straightforward.
-This is because $\tau_{\rm a}/\tau_{\rm e}$ depends on the aspect ratio of the disk but not the surface density.
-The disk thermal structure relies on the heating mechanism and therefore a specific disk model.
-As such, our framework cannot predict whether the planet pairs in the stable or overstable regimes will remain trapped as gas disk dissipates.
-In addition, the disk is assumed to be laminar here.
-In a more realistic turbulent disk, density fluctuations generated by turbulence can result in additional stochastic torques (Goldberg & Batygin, 2023; Rein, 2012).
-This effect cannot be captured by our simplified approach that assumes static migration and eccentricity damping timescales.
-However, the outcomes of resonance trapping may be altered by considering the stochastic torques, which is a key focus of our future work.
+This is because $\tau_{\rm m}/\tau_{e}$ depends on the aspect ratio of the disk but not the surface density.
+The disk thermal structure relies on the heating mechanism.
+While our framework remains applicable and can predict the evolution of resonance configurations as long as $\tau_{m}$ and $\tau_{e}$ are specified, understanding how these timescales change with disk dissipation is essential for making such predictions.
+However, this aspect is not within the scope of our current study.
 5 Conclusion
-We have built an analytical framework suited to study the resonant trapping, stability, and escape conditions for a two-planet pair that undergoes the convergent type I migration in a gaseous disk.
-The $2$:$1$ first order mean motion resonance is specifically investigated and the key outcomes are illustrated in Fig 6.
-Planets can be trapped in resonances under the circumstances of slow migration and weak damping (Eqs.
-(8) and (11)).
-When the mass of inner planet is comparable or less massive than the outer planet ($q{\lesssim}1$), after capture the stability weakens as $\tau_{\rm a}/\tau_{\rm e}$ decreases.
+We have constructed an analytical framework suited to study the resonant trapping, stability, and escape conditions for two-planet systems that undergo convergent type I migration in the protoplanetary disks.
+The planet-disk interaction is assumed to be parameterized by two key quantities: angular momentum damping and eccentricity damping timescales $\tau_{\rm m}$ and $\tau_{e}$.
+We present the analytical criteria for general $j$:j1j-1 resonances and for planet pairs with arbitrary mass ratios.
+The outcomes can be classified into four regimes: direct resonance crossing, stable resonance trap, overstable resonance trap and escape after temporary trap.
+The proposed criteria and regimes, which are validated through numerical N-body simulations, can be illustrated in a $\tau_{\rm m}/\tau_{\rm e}$-$\tau_{e}$ plot (Figure 6).
+We find that the planets can be trapped in resonances under slow migration and weak damping (Eqs.
+8 and 11).
+When the mass of the inner planet is comparable to or less massive than the outer planet ($q{\lesssim}1$), after capture, the stability weakens as $\tau_{\rm m}/\tau_{e}$ decreases.
 The system transitions from a stable trap to an overstable trap (Eq.
@@
 This corresponds to the resonant states changing from statically stable to dynamically stable, and ultimately to dynamically unstable.
-In contrast, when the mass of inner planet is much more massive than the outer planet ($q{\gg}1$), the traps remain permanently stable.
-We provide the general analytical criteria for planets with an arbitrary mass ratio $q$ in Sec. 2. The corresponding formulas are also summarized as follows.
-$\begin{cases}{\displaystyle(\tau_{e}\tau_{a})_{\rm weak,e}=\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vspace{0.1cm}\\ {\displaystyle(\tau_{a})_{\rm slow,mig}=\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}],\end{cases}$ (33)
-where
-$h(q)$ $=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (34)
-$f(q)$ $=$ $1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (35)
-We validate the theoretical predictions with numerical N-body simulations (Figures 8, 8 and 9).
-Furthermore, within the framework, we show that resonance trapping cannot occur when planets migrate divergently.
-We also explored the dependence of the semi-major axis and the stellar mass based on minimum mass solar nebular (MMSN) model, which indicates planets are more likely to be resonance-trapped with increasing semi-major axis and decreasing stellar mass.
+In contrast, when the mass of the inner planet is much more massive than the outer planet ($q{\gg}1$), the traps remain permanently stable.
+The key formulas of different regimes are presented in Eqs.
+(13) and (24).
+Future studies are needed to connect this theoretical framework with observations (Dai et al., 2024; Hu et al., 2025).
 Acknowledgements.
-This work is supported by REBOUND (Rein & Liu, 2012) and REBOUNDX (Tamayo et al., 2020)
+The authors appreciate the constructive comments from the referee.
+BL is supported by the National Key R&D Program of China (2024YFA1611803), the National Natural Science Foundation of China (Nos.
+12222303 and 12173035), and the start-up grant of the Bairen program from Zhejiang University.
+The simulations and analysis presented in this article were carried out on the SilkRiver Supercomputer of Zhejiang University.
+Software: REBOUND (Rein & Liu, 2012) and REBOUNDX (Tamayo et al., 2020).
 Appendix A: The derivation of the excited eccentricity in the classical pendulum model
@@
 where the contribution from $\dot{\varpi_{\rm i}}$ and $\dot{\varpi_{\rm o}}$ are neglected.
-With no additional effects from the gaseous disk such as orbital decay and eccentricity damping, Lagrange’s equations can be written as
+With no additional effects from the gaseous disk such as angular momentum damping and eccentricity damping, Lagrange’s equations can be written as
 $\dot{n_{\rm i}}$ $=-3(j-1)\mu_{\rm o}n_{\rm i}^{2}\alpha(e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}}),$ (A2a)
@@
 $\ddot{\varphi}{=}3[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}](e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}})\sin{\varphi},$ (A3)
-$\dot{e_{\rm i}}f_{\rm d,i}-\dot{e_{\rm o}}f_{\rm d,o}^{{}^{\prime}}=-(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}})\sin{\varphi},$ (A4)
+$\dot{e_{\rm i}}f_{\rm d,i}-\dot{e_{\rm o}}f_{\rm d,o}^{{}^{\prime}}=-(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})\sin{\varphi},$ (A4)
 where we assume $\sin{\varphi}\equiv\sin{\varphi_{\rm i}}=-\sin{\varphi_{\rm o}}$. The equation for $\ddot{\varphi}$ can be simplified to the form
@@
 We can relate the variation of $\dot{\varphi}$ to the variation in $e_{\rm i,o}$ by means of Eq. (A4). This gives
-$d(e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{d,o^{{}^{\prime}}}){=}\pm\frac{(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}})}{\omega_{0}}\sin{\frac{\varphi}{2}}d{\varphi}.$ (A8)
+$d(e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{d,o^{{}^{\prime}}}){=}\pm\frac{(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})}{\omega_{0}}\sin{\frac{\varphi}{2}}d{\varphi}.$ (A8)
 Substitute $\omega_{0}$ into Eq. (A8) and integrate it, we obtain
 $e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{\rm d,o}^{{}^{\prime}}=-\frac{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}\cos{\frac{\varphi}{2}}.$ (A9)
-The maximum eccentricity occurs when $\varphi=0$, which provides the eccentricity described in Sec. 2.1.1.
+The maximum eccentricity occurs when $\varphi=0$, which provides the eccentricity described in Section 2.1.1.
 Appendix B: Extension of stability and escape criteria to arbitrary mass ratios
@@
 For the cases of $q\ll 1$ and $q\gg 1$, the system follows the same Hamiltonian equation:
-$-H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (B1)
+$H=\eta\Phi+\Phi^{2}-\Phi^{1/2}\cos{\varphi},$ (B1)
 where
-$\begin{cases}{\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}]^{2/3}\frac{e_{\rm i}^{2}}{4},\ \ \hfill[\mbox{$q\ll 1$}],\vspace{0.1cm}}\\ {\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm i}}\right)^{2/3}[\frac{j^{2}}{f_{\rm d,o}^{{}^{\prime}}}]^{2/3}\frac{e_{\rm o}^{2}}{4},\ \ \hfill[\mbox{$q\gg 1$}].}\end{cases}$ (B2)
+$\begin{cases}{\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}]^{2/3}\frac{e_{\rm i}^{2}}{4},\ \ \hfill[\mbox{$q\ll 1$}],\vskip 2.84544pt}\\ {\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm i}}\right)^{2/3}[\frac{j^{2}}{f_{\rm d,o}^{{}^{\prime}}}]^{2/3}\frac{e_{\rm o}^{2}}{4},\ \ \hfill[\mbox{$q\gg 1$}].}\end{cases}$ (B2)
 Therefore, we hypothesize the existence of a form for $\Phi$ that satisfies the Hamiltonian equation and degenerates to Eq. (B3) in the extreme mass ratio cases:
@@
 where $e^{2}=e_{\rm i}^{2}+e_{\rm o}^{2}$.
-When $\eta_{\rm eq}{<}-3/2$ (equivalently $\Phi_{\rm eq}>1$), a separatrix point occurs and the system can cross this separatrix and escape from resonance (see Sec.
-2.2.2).
+When $\eta_{\rm eq}{<}-3/2$ (equivalently $\Phi_{\rm eq}>1$), a separatrix point occurs and the system can cross this separatrix and escape from resonance (see Section 2.2.2).
 By substituting $e_{\rm i,eq}$ and $e_{\rm o,eq}$ (Eq.
@@
 (B3), we obtain the general escape criterion as
-$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}.$ (B4)
+$\frac{\tau_{\rm m}}{\tau_{e}}<\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}.$ (B4)
 where
-$h(q)$ $=$ $\tau_{a}/\tau_{e}\left(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2}\right)$ (B5)
-$=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}}.$
+$h(q)$ $=$ $\tau_{\rm m}/\tau_{e}\left(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2}\right)$ (B5)
+$=$ $\frac{1}{(1+q\sqrt{\alpha})}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}}.$
 B.2: Stability criterion
@@
 When the outer planet is comparable or more massive than the inner planet ($q\lesssim 1$), the system may become overstable (see Eq.
-22).However, all traps remain stable when the inner planet is more massive.
+22).
+However, all traps remain stable when the inner planet is more massive.
 Therefore, we introduce a factor $f(q)$ to fit the general stability criterion, which is given by
-$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q),$ (B6)
+$\frac{\tau_{\rm m}}{\tau_{e}}>\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q),$ (B6)
 where
@@
 When $e_{\rm o,eq}>e_{\rm i,eq}$, $f(q)<0$, indicating that the system is dominated by the inner planet and all traps are stable traps.
-B.3: Dependence on the mass ratio $q$
-We run simulations to testify to our proposed overstability and escape criteria.
-Figure B.1 shows the numerical results of $\tau_{\rm a}/\tau_{\rm e}$ as a function of $q$.
-We fix $m_{\rm i}{=}10\ M_{\oplus}$ and set $\tau_{\rm a}{=}3\ (\tau_{\rm a})_{\rm slow,mig}$ to fulfill the trapping criteria.
-The solid lines denote the analytical criteria, while the colored dots represent the simulation results.
-The numerical results agree pretty well with the analytical expectation at $q{\leq}1$.
-The simulations depict a transition from stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes as $\tau_{\rm a}/\tau_{\rm e}$ decreases, and our analytical criteria separate these three regimes.
-In contrast, when $q\geq 1$, there are still two points in Figure B.1 where the simulations do not fully match the analysis.
-First, the stability/overstability and escape criteria exhibit the opposite dependence $q$, with the two lines intersecting at $q{\sim}2$.
-For $q{\geq}2$, one would expect that the systems reside region below the overstability criterion (yellow line) correspond to escape.
+B.3: Dependence on planet mass ratio
+We run simulations to test our proposed overstability and escape criteria with different planet mass ratios by fixing $m_{\rm i}{=}10\ M_{\oplus}$.
+We set $\tau_{\rm m}{=}3\ (\tau_{\rm m})_{\rm slow,mig}$ to fulfill the trapping criteria and vary $\tau_{e}$.
+The result is presented in Figure B.1, where the solid and dashed lines denote the analytical criteria from ours and from Deck & Batygin (2015).
+The simulations agree pretty well with the analytical expectation at $q{\leq}1$, with our criteria slightly better than Deck & Batygin (2015).
+The simulations depict a transition from stable trap (green squares) to overstable trap (yellow circles), and eventually become escape (purple triangles) as $\tau_{\rm m}/\tau_{e}$ decreases, and our analytical criteria separate these three regimes.
+However, at $q{\geq}1$, there are still two mismatches between simulations and analysis in Figure B.1.
+First, the stability/overstability and escape criteria exhibit an opposite trend in q, with the two lines intersecting at $q{\sim}2$.
+For $q{\gtrsim}2$, one would expect that the systems in the region below the overstability criterion (yellow line) should escape.
 However, simulations show that all systems in this region are trapped overstably (no escape occurs).
-We attribute this to the fact that when $q{\geq}2$, the angular momentum deficit (AMD) (Laskar 1997) of the inner planet is larger than that of the outer planet, limiting the amplitude of oscillations in overstability (see Appendix B.4 for detailed derivation).
-As a result, the perturbations cannot grow sufficiently to cause escape from resonance.
-The dashed purple line in Figure B.1 represents the AMD criterion when the second-order terms of the eccentricities are considered.
-This new criterion is in good agreement with our numerical simulations.
-Second, our overstable criterion suggests $\tau_{\rm a}/\tau_{\rm e}{<}0$ at $q{\geq}3$, indicating that all traps should be stable in this regime.
-However, in Figure B.1, it is evident that near $q\sim 4$ and $\tau_{a}/\tau_{e}\sim 10^{2}$, the planets remain in an overstability state.
-This is because we only considered the first-order term of the disturbing function.
-Second-order terms are neglected here, weakening the inner planet effect on the outer planet and leading to an overstable trap.
-Figure B.1: Simulation results of $2:1$ MMR stability in $q-\tau_{a}/\tau_{e}$ parameter space with the mass of the inner planet $m_{\rm i}=10\ \times M_{\oplus}$.
-Green points indicate stable resonance capture, yellow points indicate overstable resonance capture, and purple points indicate the planets eventually escape from the resonance.
-The lines represents our analytical criteria.
-It can be observed that our criteria overall align well with the numerical results.
+We attribute this to the fact that when $q{\gtrsim}2$, the angular momentum deficit (AMD) (Laskar 1997) of the inner planet is higher than that of the outer planet.
+The exchange of angular momentum between these two planets weakens the oscillations of the amplitude of the outer planet’s eccentricity (see Appendix B.4 for detailed derivation).
+Thus, the perturbations cannot grow sufficiently to cause escape, and the system is refined in the overstable trap regime.
+The solid blue line in Figure B.1 represents the AMD criterion up to the second-order eccentricity terms.
+For comparison, the blue dotted line is the AMD criterion with only the first-order eccentricity terms.
+We find that the the AMD criterion with the second-order eccentricity terms separates the overstable trap and escape regimes at $q{\gtrsim}2$, which is in good agreement with our simulations.
+Second, our overstable criterion suggests at $q{\geq}3$, all traps are stable.
+However, in Figure B.1, it is evident that near $q\sim 4$ and $\tau_{\rm m}/\tau_{e}\sim 2\times 10^{2}$, the planets remain in an overstable state.
+We suspect the reason might be that we only considered the first-order term of the disturbing function.
+When the second-order term is included, the angular momentum transfer from the inner planet to the outer one becomes weaker.
+This leads to the shift of the regimes between stable and overstable to a higher $q$.
+A detailed analytical investigation for the offset seen in Fig.
+B.1 is nevertheless beyond the scope of this work.
+Besides, we also extend the test to the 33:22 MMR in Fig. B.2. We also find that our analysis agrees well with the numerical simulations.
+Figure B.1: Simulations of 22:11 MMR capture and stability in $q-\tau_{\rm m}/\tau_{e}$ parameter space.
+The mass of the inner planet $m_{\rm i}=10\ M_{\oplus}$.
+Green: stable resonance capture.
+Yellow: overstable resonant capture.
+Purple: resonance escape after temporary capture.
+The solid lines are the criteria derived in this study, whereas the dashed lines represent the stability and escape criteria of Deck & Batygin (2015) (their Eqs.
+15 and 21).
+The blue solid and dotted lines represent our second-order (Eq.
+B16) and first-order (Eq.
+B17) AMD criteria, respectively.
+Figure B.2: Same as Fig B.1, but for the 33:22 MMR.
 B.4: Second-Order AMD criterion
-We expand the disturbing functions to the second order in $e_{\rm i}$ and $e_{\rm o}$, which gives
+We expand the disturbing functions up to the second order in $e_{\rm i}$ and $e_{\rm o}$, which gives
 $R_{\rm D}^{\rm sec}$ $=$ $K_{1}(e_{\rm i}^{2}+e_{\rm o}^{2})+K_{2}e_{\rm i}e_{\rm o}\cos{(\varpi_{\rm i}-\varpi_{\rm o})},$ (B8)
@@
 The angular momentum deficit (AMD) (Laskar 1997) is given by
-$AMD\sim\mu\sqrt{a}(1-\sqrt{1-e^{2}})\sim\mu\sqrt{a}e^{2}.$ (B14)
+$AMD_{\rm i,o}\sim\mu_{i,o}\sqrt{a_{\rm i,o}}(1-\sqrt{1-e_{\rm i,o}^{2}})\sim\mu_{\rm i,o}\sqrt{a_{\rm i,o}}e_{\rm i,o}^{2}.$ (B14)
 When the AMD of the inner planet exceeds that of the outer planet, the condition
@@
 (B12) and (B13) to eliminate either $e_{\rm i}$ and $e_{\rm o}$ under the condition $\dot{\varpi_{\rm i}}=\dot{\varpi_{\rm o}}$, we obtain the relationship between $e_{\rm i,o}$ and $q$.
-Assuming $e_{\rm i}$ takes the form: $e_{\rm i}^{2}={\tau_{e}}/{2j\tau_{a}}$, and substituting this into the derived form, we obtain our AMD criterion as
-$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm AMD}=\frac{2}{j}\left[\frac{(f_{3}-K_{1})+q^{{}^{\prime}2}(f_{5}+K_{1})}{q^{{}^{\prime}}f_{\rm d,o}^{{}^{\prime}}+f_{\rm d,i}}\right]^{2},$ (B16)
-where $q^{{}^{\prime}2}\equiv q\sqrt{\alpha}$. The AMD criterion aligns well with the numerical results (Figure B.1).
+Assuming $e_{\rm i}$ takes the form: $e_{\rm i}^{2}={\tau_{e}}/{j\tau_{\rm m}}$, and substituting this into the derived form, we obtain our second-order AMD criterion as
+$\left(\frac{\tau_{\rm m}}{\tau_{e}}\right)<\frac{1}{j}\left[\frac{(f_{3}-K_{1})+q^{{}^{\prime}2}(f_{5}+K_{1})}{q^{{}^{\prime}}f_{\rm d,o}^{{}^{\prime}}+f_{\rm d,i}}\right]^{2},$ (B16)
+where $q^{{}^{\prime}2}\equiv q\sqrt{\alpha}$. We can also obtain the first-order AMD criterion by substituting Eq. (5) and (4b) into Eq. (B15):
+$q>\left(\frac{f_{\rm d,i}}{f_{\rm d,o}^{{}^{\prime}}}\right)^{2}\frac{1}{\sqrt{\alpha}}.$ (B17)
+By comparison, the inclusion of the second-order terms induce a $\tau_{\rm m}/\tau_{e}$ dependency, which is absent in the first-order approximation.
+We plot our AMD criteria in Figs.
+B.1 and B.2 and find that the second-order AMD criterion exhibits improved agreement with the numerical results.
+C: Extension to Unequal Eccentricity Damping Timescales
+We extend the framework to the planet pair systems with non-equal eccentricity damping. Equation (3) can be written as
+$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})$ (B18)
+$=\frac{1}{\tau_{\rm m}}+\frac{e_{\rm o}^{2}}{\tau_{e\rm,o}}-\frac{e_{\rm i}^{2}}{\tau_{e\rm,i}}.$
+The equilibrium eccentricities read
+$e_{\rm i,eq}^{2}$ $=$ $\frac{\tau_{e,\rm i}}{\tau_{\rm m}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}}{jf_{\rm d,i}^{2}+(j-1)\frac{\tau_{e\rm,i}}{\tau_{e\rm,o}}f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (B19a)
+$e_{\rm o,eq}^{2}$ $=$ $\frac{\tau_{e,\rm i}}{\tau_{\rm m}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)\frac{\tau_{e\rm,i}}{\tau_{e\rm,o}}f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}}.$ (B19b)
+The key criteria can be re-formulated as
+$\begin{cases}{\displaystyle\tau_{e,\rm i}\tau_{\rm m}>\frac{1}{(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)\frac{\tau_{e,\rm i}}{\tau_{e,\rm o}}f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vskip 2.84544pt\\ {\displaystyle\tau_{\rm m}>\frac{1}{[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}],\vskip 2.84544pt\\ {\displaystyle\frac{\tau_{\rm m}}{\tau_{e,\rm i}}>\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vskip 2.84544pt\\ {\displaystyle\frac{\tau_{\rm m}}{\tau_{e,\rm i}}<\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}],\end{cases}$ (B20)
+where
+$h(q)$ $=$ $\frac{1}{(1+q\sqrt{\alpha})}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)\frac{\tau_{e,\rm i}}{\tau_{e,\rm o}}f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (B21)
+$f(q)$ $=$ $1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (B22)
+These criteria have also been validated and compared with those of Deck & Batygin (2015) through simulations by incorporating different eccentricity dampings of two planets at $\tau_{e,\rm i}/\tau_{e,\rm o}{=}2$ (see Fig.
+C.1).
+However, we also find that the analytical formulas increasingly deviate from the simulation results as the ratio $\tau_{e,i}/\tau_{e,o}$ increases, particularly for values greater than 10.
+We suspect that this is because, in our derivation, $\cos{\varphi_{\rm i}}{=}\cos{\varphi_{\rm o}}{=}1$ is assumed, whereas it can not hold for extreme eccentricity damping ratios.
+A detailed investigation is nevertheless beyond the scope of this work.
+Figure C.1: Same as Fig 9, but $\tau_{e,\rm i}/\tau_{e,\rm o}=2$.